### v1

# Certifiably Optimal Anisotropic Rotation Averaging

Abstract
Rotation averaging is a key subproblem in applications of computer vision and robotics. Many methods for solving this problem exist, and there are also several theoretical results analyzing difficulty and optimality. However, one aspect that most of these have in common is a focus on the isotropic setting, where the intrinsic uncertainties in the measurements are not fully incorporated into the resulting optimization task. Recent empirical results suggest that moving to an anisotropic framework, where these uncertainties are explicitly included, can result in an improvement of solution quality. However, global optimization for rotation averaging has remained a challenge in this scenario. In this paper we show how anisotropic costs can be incorporated in certifiably optimal rotation averaging. We also demonstrate how existing solvers, designed for isotropic situations, fail in the anisotropic setting. Finally, we propose a stronger relaxation and show empirically that it is able to recover global optima in all tested datasets and leads to a more accurate reconstruction in all but one of the scenes.
1 Introduction
Rotation averaging has been a topic of interest now for some time, dating back more than two decades, with early work such as [18]. Govindu’s groundbreaking work has since been followed up by many others and the study of rotation averaging is now a very important area.
In large part this is because of the great importance it has for many problems within computer vision and robotics. Traditionally, it has been a core component of non-sequential structure from motion (SfM), see e.g. [26, 27, 32]. The key strength here is that rotation averaging optimizes the orientations of all the cameras at the same time, allowing recovery methods to avoid incrementally increasing errors (drift) [12] that easily occur in sequential SfM methods. More recently, [5, 7, 8] also point out the importance of rotation averaging for improving accuracy and speed of simultaneous localization and mapping (SLAM).
The main difficulty in solving the rotation averaging problem is that rotations reside in a nonlinear manifold $SO(3)$. Some approaches [13, 20, 19, 9, 22, 39] address this challenge by using local optimization over $SO(3)^{n}$ via explicit parametrizations. These, and other methods of local optimization all work by moving along descent directions, which may lead to a global optima in practical and sufficiently “nice” situations. Wilson et al. [44, 43] study when this occurs, and when local optimization is “hard” due to the existence of “bad” local minima. They conclude that this depends on both the measurement noise and on the algebraic connectivity of the graph representing the problem.
More recent methods for solving rotation averaging problems formulate a non-convex quadratically constrained quadratic problem (QCQP) and relax it to a semidefinite program (SDP). The quadratic constraints force the desired matrices to be orthogonal. The non-quadratic determinant constraint is often dropped, effectively replacing optimization over $SO(3)$ with one over $O(3)$ [4, 14, 33, 25, 36, 10]. The SDP formulation is a relaxation of the QCQP and results in a strictly less or equal optimal value than the QCQP. There are, however, several theoretical results [16, 20, 36] that point to this “gap” often being non-existent, giving a tight lower bound allowing the recovery of a “certifiably correct” solution. In [16] an explicit bound depending on the measurement noise and algebraic connectivity of the graph is given, in analogy to the local optimization setting.
While the above methods and algorithms are often certifiably correct, they minimize the isotropic chordal distance and, therefore, ignore the relative rotation uncertainties. The anisotropic version of the chordal distance allows to explicitly include uncertainties obtained by local two-view optimization into the rotation averaging framework. Recent empirical results such as [46] suggest that accounting for uncertainties in the optimization could “largely improve reconstruction quality”. Therefore the isotropic assumption discussed above can be seen as a disadvantage.
In this paper, we develop a certifiably optimal SDP-formulation able to optimize the anisotropic chordal distance. In light of the above, this is an improvement of many earlier results. However, global optimization will become more challenging in this case, as—based on our analysis—a direct modification of the objective function results in an SDP that rarely provides a tight lower bound. The main reason for this is, that optimization of an anisotropic cost over $O(3)$ typically results in a solution outside $SO(3)$, implying that the determinant constraint on rotation matrices cannot be dropped in order to expect tight semidefinite relaxations. As mitigation we present a new relaxation that is capable of further constraining the solution to the convex hull of $SO(3)$, $\operatorname{conv}(SO(3))$.
Further, we empirically verify—on a number of synthetic and real datasets—that this new relaxation is sufficient to recover the globally optimal solution, and that our anisotropic model generally achieves solutions that are more accurate than the standard isotropic chordal penalty.
In summary, our main contributions are:
  • We show how anisotropic costs can be incorporated in certifiably correct rotation averaging.
  • We provide an analysis of the objective function that explains why regular solvers only enforcing $O(3)$ membership usually fail in the anisotropic setting.
  • We present a stronger convex relaxation able to enforce $\operatorname{conv}(SO(3))$ membership and verify empirically, that the proposed formulation is able to recover a global optimum in all tested instances. To our knowledge, this is the first formulation to yield tight relaxations.
2 Related Work
There are many ways of formulating the rotation averaging problem and representing rotations; [3, 30] represent the rotations using vectors and optimize over them, while [17] uses quaternions. Quaternions have some advantages, but the fact that unit quaternions form a double cover of $SO(3)$ means that quaternion distances are less straightforward and a second optimization step is needed to determine their appropriate signs. In this paper we will only consider matrix representations.
The works [2, 21] also considers certifiable optimization of the anisotropic chordal distance. However, their SDP relaxation is based on the Cayley mapping and the relation to our formulation is somewhat unclear. In addition, [2, 21] have to introduce redundant constraints in order to get a “reasonably tight” SDP relaxation. It is likely, in the light of our analysis in this paper, and the discussion in [3] that this is a result of the authors only enforcing $O(3)$ membership for their optimization.
It is also worth mentioning that—while this paper looks at the anisotropic $L_{2}$ chordal distance—, it is possible to consider other distances between rotations as well as their robust counterparts. E.g., [43, 44] consider the geodesic distance (i.e. the one on the manifold of $SO(3)$). Further, in [9, 39, 42] the authors analyze and look at different (robustified) distances in order to obtain a “robust” formulation (resulting in the solution being less susceptible to outliers). Outliers are handled via the introduction of auxiliary variables in the SDP formulation in [35, 6].
From an algorithmic point of view, the SPD relaxation can be solved by general-purpose solvers, e.g. [1]. There are also several dedicated solvers, constructed specifically for the rotation averaging problem. [16, 33] use coordinate descent (fixing all but one row/column and minimizing), [25] use a primal-dual update rule, and in [10, 45] the authors combine a global coordinate descent approach with a local step. An early approach [40] looks at eigenvectors of a specific matrix. A particularly powerful method used in [4, 14] is that of the Riemannian staircase, where one optimizes using block matrices of a certain dimension and increases (“climb the staircase”) the dimension until a global optimum is found. The authors show that such a method can be made highly effective.
3 Certifiably Optimal Rotation Averaging
In the context of structure from motion and SLAM, the goal of rotation averaging is to estimate absolute camera orientations from relative rotation measurements between pairs of calibrated cameras [18, 20, 17, 40]. If $R_{i}$ and $R_{j}$ are rotation matrices, encoding the orientation of cameras $P_{i}=\begin{bmatrix}R_{i}&t_{i}\end{bmatrix}$ and $P_{j}=\begin{bmatrix}R_{j}&t_{j}\end{bmatrix}$, the relative rotation $R_{ij}$ between the two cameras, that is, $P_{i}$’s orientation in the camera coordinate system of $P_{j}$ is $R_{ij}=R_{i}R_{j}^{\top}$. Given estimates $\tilde{R}_{ij}$ of $R_{ij}$, typically obtained by solving two view relative pose [28], the traditional rotation averaging formulation [16, 14, 33] accounts for estimation error using the chordal distance $\|\tilde{R}_{ij}-R_{i}R_{j}^{\top}\|_{F}$ and aims to solve
$\min_{R_{i}\in SO(3)}\sum_{i\neq j}\|\tilde{R}_{ij}-R_{i}R_{j}^{\top}\|_{F}^{2}.$ (1)
The derivation of a strong convex relaxation relies on two observations: Firstly, since the set of rotation matrices have constant (Frobenius) norm, the objective function can be replaced by $-\sum_{i\neq j}\langle\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle$. If we let $\mathbf{R}=\begin{bmatrix}R_{1}^{\top}&R_{2}^{\top}&\ldots\end{bmatrix}^{\top}$ and $\mathbf{\tilde{R}}$ be the matrix containing the blocks $\tilde{R}_{ij}$ (and zeros where no relative rotation estimate is available) we can simplify the objective to $-\langle\mathbf{\tilde{R}},\mathbf{R}\mathbf{R}^{\top}\rangle$, which is linear in the relative rotations. Secondly, by orthogonality we have $R_{i}R_{i}^{\top}={\mathtt{I}}$, resulting in the Quadratically Constrained Quadratic Program (QCQP)
$p^{*}:=\min_{\mathbf{R}}\,$ $-\langle\mathbf{\tilde{R}},\mathbf{R}\mathbf{R}^{\top}\rangle\quad\text{s.t. }\forall i:R_{i}R_{i}^{\top}={\mathtt{I}}.$ (2)
QCQP are typically non-convex but can be relaxed to convex linear semidefinite programs (SDP) by replacing $\mathbf{R}\mathbf{R}^{\top}$ with a positive semidefinite matrix $\mathbf{X}$ (technically by taking the Lagrange dual twice), yielding the relaxation
[…]
The above program is convex and can therefore be solved reliably using general purpose [1, 23] or specialized solvers [14, 4, 16]. The difference $p^{*}-d_{2}^{*}$ between objective values is referred to as the duality gap. It is clear that we can construct $\mathbf{X}=\mathbf{R}\mathbf{R}^{\top}$ from a solution $\mathbf{R}$ of (2), which is feasible in (SDP-O(3)-iso) and therefore $p^{*}\geq d_{2}^{*}$. Moreover, if the solution $\mathbf{X}$ to (SDP-O(3)-iso) has $\operatorname{rank}(\mathbf{X})=3$ then it can be factorized into $\mathbf{X}=\mathbf{R}\mathbf{R}^{\top}$, with $\mathbf{R}$ being feasible in (2), which yields $p^{*}=d_{2}^{*}$ and certifies that $\mathbf{R}$ is globally optimal in (2). Thus, a solution to the original non-convex QCQP is obtained through optimization via convex SDP program.
We remark that the use of a linear objective is important to obtain a strong relaxation. Linear objectives admit solutions at extreme points (on the boundary) of the feasible set which are often of low rank. It has been shown, both theoretically and empirically, [16, 35, 40, 8, 36] that the duality gap of synchronization problems is often zero under various bounded noise regimes. In [16], explicit bounds depending on the algebraic connectivity of the camera graph are given. For example, for fully connected graphs the duality gap can be shown to be zero if there is a solution with an angular error of at most $42.9^{\circ}$—which is a rather generous measurement error.
4 Anisotropic Rotation Averaging
While the use of the chordal distance often results in tight SDP relaxations, a downside is that it does not take into account the uncertainty of the relative rotation estimates $\tilde{R}_{ij}$. For two-view relative pose problems, it is frequently the case that the reprojection errors are much less affected by rotations in certain directions than others. In Figure 1 we illustrate this on a real two-view problem. The left image shows the ground truth solution (obtained after bundle adjustment). To generate the graphs to the right, we sampled rotations of the second camera around its x-,y-, and z-axes between $-5$ and $5$ degrees. For each sampled rotation, we found the best 3D point locations and position of the second camera, and computed the sum-of-squared (calibrated) reprojection errors (solid curves). The result shows that rotations around the black/upward axis affects the reprojection errors significantly less than the others, resulting in a direction of larger uncertainty. By performing the sensitivity analysis (i.e. by looking at the singular values of the Jacobian) one can arrive at the similar conclusion. It is therefore desirable to propagate this information to the rotation averaging stage to increase the accuracy of the model. The dashed curves (that are almost indistinguishable from the solid ones) show the approximations, presented below, that can be used for this purpose. In the supplementary material, we also show how isotropic rotation averaging gets negatively affected by the single noisy relative rotation (while the proposed method presented further remains unaffected).
Figure 1: Ground truth solution to two view problem (left). Sum-of-squared reprojection errors (right, solid curves) obtained when rotating the second camera around the black, green and red axes (from -5 to 5 degrees) and anisotropic quadratic approximations (dashed curves) that can be used in rotation averaging.
4.1 Incorporating uncertainties
Consider a relative rotation $\tilde{R}$ that has been optimized to a minimum using a nonlinear least squares approach in a local two-view optimization. Using the exponential map we write $R=e^{[\Delta\bm{\omega}]_{\times}}\tilde{R}$ [^1] Other parameterizations are discussed in the supplementary material., where $\Delta\bm{\omega}$ is an axis-angle representation of the deviation from $\tilde{R}$, i.e. the axis-angle of $R\tilde{R}^{\top}$. Since $\tilde{R}$ is a minimum, gradient terms vanish and the two-view objective can locally be approximated by a quadratic function (ignoring an inconsequential constant)
$\tfrac{1}{2}\Delta\bm{\omega}^{\top}H\Delta\bm{\omega}.$ (3)
The positive semidefinite matrix $H=J^{\top}J$ (where $J$ is the Jacobian of the residual function) approximates the growth of the objective function in a region around the minimum $\tilde{R}$. We remark that the initial two-view objective function can depend on additional parameters such as camera positions and 3D point locations. Here we assume that these “nuisance” parameters have been marginalized out and therefore work with a reduced objective solely in terms of the rotation parameters $\Delta\bm{\omega}$ (using the Schur complement to factor out additional unknowns).
We assume Laplace’s approximation to be valid at least locally in order to quantify the uncertainty of $\Delta\bm{\omega}$, i.e. identify the precision matrix with the Hessian $H$. While in general the uncertainty of a $3\times 3$ matrix $R$ is represented by a $9\times 9$ covariance/precision matrix, we aim for a linear cost (in terms of rotation matrices) in order to achieve a strong relaxation, which means that the objective is restricted to an inner product between $R$ and some other (input-dependent) matrix. We first find a matrix $M$ such that
$\begin{split}\Delta\bm{\omega}^{\top}H\Delta\bm{\omega}&=\operatorname{tr}\left([\Delta\bm{\omega}]_{\times}^{\top}M[\Delta\bm{\omega}]_{\times}\right)\\ &=-\operatorname{tr}\left(M[\Delta\bm{\omega}]_{\times}^{2}\right).\end{split}$ (4)
Using the identity $[\Delta\bm{\omega}]_{\times}^{2}=\Delta\bm{\omega}^{\top}\Delta\bm{\omega}{\mathtt{I}}-\Delta\bm{\omega}\Delta\bm{\omega}^{\top}$, we obtain that $H=\operatorname{tr}(M){\mathtt{I}}-M$. Taking the trace on both sides further shows that $\operatorname{tr}(H)=2\operatorname{tr}(M)$, resulting in
$M=\tfrac{\operatorname{tr}(H)}{2}{\mathtt{I}}-H.$ (5)
Next, we deduce via the Taylor expansion of the exponential map that
$[\Delta\bm{\omega}]_{\times}\approx R\tilde{R}^{\top}-{\mathtt{I}},$ (6)
with equality up to first order terms. This yields the objective corresponding to (3) based on (4),
$\tfrac{1}{2}\operatorname{tr}\big{(}(R\tilde{R}^{\top}-{\mathtt{I}})^{\top}M(R\tilde{R}^{\top}-{\mathtt{I}})\big{)}.$ (7)
Similarly to the isotropic case, by leveraging the properties of rotation matrices and by omitting constants, this objective can be reduced to a linear one [^2] In the supplementary material we show that we arrive at the same objective using classical first-order uncertainty propagation.,
$-\operatorname{tr}\big{(}M\tilde{R}R^{\top}\big{)}=-\langle M\tilde{R},R\rangle.$ (8)
We observe that this objective can be viewed as a negative log-likelihood of a Langevin distribution $\mathcal{L}(R,M)$, e.g. [11], with a density function of the type
$p(\tilde{R})\propto e^{\langle M\tilde{R},R\rangle},$ (9)
and that generating rotations $\tilde{R}$ according to (9) is approximately (up to second order terms) equivalent to generating axis-angle vectors $\Delta\bm{\omega}$ from a normal distribution $\mathcal{N}(0,H^{-1})$ with density
$p(\Delta\bm{\omega})\propto e^{-\frac{1}{2}\Delta\bm{\omega}^{\top}H\Delta\bm{\omega}}.$ (10)
Hence, the above derivations suggest a way of generalizing the rotation averaging approach by replacing the relative estimates $\tilde{R}_{ij}$ with a weighted version $M_{ij}\tilde{R}_{ij}$, where $M_{ij}$ is computed from the Hessian $H_{ij}$ of the relative pose problem as outlined above. Therefore, our proposed objective to minimize is
$-\sum\nolimits_{i\neq j}\langle M_{ij}\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle=-\langle\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle,$ (11)
where $\mathbf{N}$ is a symmetric block matrix containing the blocks $(M_{ij}\tilde{R}_{ij})^{\top}$ for $i<j$, 00 when $i=j$, and $M_{ji}\tilde{R}_{ji}$ for $i>j$.
We conclude this section by remarking that—in contrast to $H$—the matrix $M$ will generally not be positive semidefinite. The following lemma characterizes its eigenvalues
Lemma 1.
If $H\succeq 0$ then $M=\frac{\operatorname{tr}(H)}{2}{\mathtt{I}}-H$ has eigenvalues $\lambda_{1}\geq\lambda_{2}\geq|\lambda_{3}|$.
Proof.
If $\eta_{1}\geq\eta_{2}\geq\eta_{3}\geq 0$ are eigenvalues of $H$ then $\frac{\operatorname{tr}(H)}{2}-\eta_{i}$, $i=1,2,3$ are the eigenvalues of $M$. Sorted in decreasing order we get $\lambda_{1}=\frac{1}{2}(\eta_{1}+\eta_{2}-\eta_{3})$, $\lambda_{2}=\frac{1}{2}(\eta_{1}-\eta_{2}+\eta_{3})$ and $\lambda_{3}=\frac{1}{2}(-\eta_{1}+\eta_{2}+\eta_{3})$. Only $\lambda_{3}$ can be negative and clearly $\lambda_{2}\pm\lambda_{3}\geq 0$. ∎
The case $\lambda_{3}<0$ occurs when the leading eigenvalue of $H$ is significantly larger than the other two, i.e. the estimate $\bm{\omega}$ is more certain in a specific direction. This is actually a rather common scenario in image-based estimation as the in-plane part of the rotation is usually better constrained than the out-of-plane rotation angles.
4.2 Global solutions: $O(3)$ vs. $SO(3)$
Incorporating the anisotropic objective (11) in the standard relaxation (SDP-O(3)-iso) may seem like a straightforward extension. We refer to this approach as SDP-O(3)-aniso. The source of the problem with this approach is that it ignores the determinant constraint and optimizes over $O(3)$. To illustrate this problem, we ran $1000$ synthetic anisotropic problem instances (the data generation protocol follows Section 5.1 with the eigenvalues of inverse Hessians sampled from $[0.1,1]$). Fig. 2 shows that the standard rotation averaging relaxation is never able to recover a rank-$3$ solution [^3] To determine rank, we used the smallest number $N$ of singular values (i.e., the first $N$ largest values) that sum up to $>99.9\%$ of the total sum. for any of the problem instances. In contrast, our relaxation SDP-cSO(3), presented further in Section 4.3, returns rank-$3$ solutions in all cases.
Figure 2: The results of running 1000 synthetic instances of rotation averaging with an anisotropic objective. In contrast to the proposed approach the standard convex relaxations give no solutions of rank $3$.
To better understand the problem, we consider terms of type $-\langle M\tilde{R},R\rangle$ that constitute the objective function. As we have seen in Section 4.1, these are locally accurate and allow uncertainties of $\tilde{R}$ to be propagated to the rotation averaging stage. However, for successful global optimization one has to guarantee that there is no other matrix in the feasible set that gives a lower cost than $\tilde{R}$. The following theorem shows, that when viewed as a function over $O(3)$, $\tilde{R}$ does typically not result in the smallest objective value:
Theorem 1.
Let the eigenvalues of $M$ be such that $\lambda_{1}\geq\lambda_{2}\geq|\lambda_{3}|$ and $\lambda_{3}<0$. Then the minimizer of
$f(R)=-\langle M\tilde{R},R\rangle+\langle M\tilde{R},\tilde{R}\rangle,$ (12)
over $SO(3)$ is given by $R=\tilde{R}$ (with $f(\tilde{R})=0$), but the minimizer $R^{\prime}$ over $O(3)$ satisfies $f(R^{\prime})=-2|\lambda_{3}|<0$.
Proof.
We minimize $-\langle M\tilde{R},R\rangle$ by computing KKT-points over $O(3)$. Introducing a (symmetric) Lagrange multiplier $\Lambda$ for the constraint $RR^{\top}-{\mathtt{I}}=0$ and differentiating
$L(R,\Lambda)=-\langle M\tilde{R},R\rangle+\tfrac{1}{2}\langle\Lambda,RR^{\top}-{\mathtt{I}}\rangle$ (13)
with respect to $R$, gives $\tfrac{\partial L}{\partial R}=-M\tilde{R}+\Lambda R\stackrel{{\scriptstyle!}}{{=}}0$. Hence the task is to find a symmetric multiplier matrix $\Lambda$ and a rotation matrix $Q=\tilde{R}R^{\top}$ such that $MQ=\Lambda$. Since $M$ is symmetric, it has an eigen-decomposition $M=UDU^{\top}$, and we must have $(MQ)(MQ)^{\top}=MM^{\top}=UD^{2}U^{\top}=\Lambda\Lambda^{\top}$. Therefore, $\Lambda$ has the same eigenvalues as $M$ up to their signs, and consequently all solutions for $\Lambda$ are given by $\Lambda=UDSU^{\top}$, where $S=\operatorname{diag}(\pm 1,\pm 1,\pm 1)$. Further
$R=\Lambda^{-1}M\tilde{R}=UD^{-1}S^{-1}U^{\top}UDU^{\top}\tilde{R}=USU^{\top}\tilde{R}$ (14)
with the corresponding objective value
$-\langle UDU^{\top}\tilde{R},USU^{\top}\tilde{R}\rangle=\pm\lambda_{1}\pm\lambda_{2}\pm\lambda_{3},$ (15)
where $\lambda_{1},\lambda_{2},\lambda_{3}$ are the eigenvalues of $M$. According to Lemma 1 the smallest eigenvalue $\lambda_{3}$ of $M$ is negative if $M$ is indefinite. In this case the objective function is minimized over $O(3)$ by choosing $S=\operatorname{diag}(1,1,-1)$ which yields the solution $R=USU^{\top}\tilde{R}$ and $-\langle M\tilde{R},R\rangle=-\lambda_{1}-\lambda_{2}-|\lambda_{3}|$. In contrast $-\langle M\tilde{R},\tilde{R}\rangle=-\lambda_{1}-\lambda_{2}+|\lambda_{3}|$, showing that the smallest value of $f$ is $f(R)=-2|\lambda_{3}|$.
Since $SO(3)\subset O(3)$ the KKT points of $SO(3)$ are those which satisfy $\det(R)=1$, which implies that the matrix $S$ has to have either zero or two elements that are $-1$. Therefore the above choice of $S$ is infeasible, and the optimal $S$ is instead given by $S=I$, which shows that the minimum value of $f$ is $f(\tilde{R})=0$. ∎
The above result shows that loss-terms of the type $-\langle M\tilde{R},R\rangle$ will favor incorrect solutions (that can be far from the estimation $\tilde{R}$) when optimizing over $O(3)$, which explains the results in Figure 2. While strictly enforcing $R\in SO(3)$ is undesirable, as it results in non-convex constraints, the following result shows that it is enough to use $R\in\operatorname{conv}(SO(3))$ to resolve this issue.
Corollary 1.
If the eigenvalues of $M$ fulfill $\lambda_{1}\geq\lambda_{2}\geq|\lambda_{3}|$ then the function (12) is non-negative on $\operatorname{conv}(SO(3))$.
Proof.
For a linear objective function it is well known that the optimal value is attained in an extreme point. For any set $C$, the extreme points of $\operatorname{conv}(C)$ are always contained in $C$ [34]. Further, since for every point $Q\in SO(3)$ the linear function $-\langle Q,R\rangle$ has its unique minimum in $R=Q$, it can be deduced that the set of extreme points of $\operatorname{conv}(SO(3))$ is exactly $SO(3)$. Thus, the minimum value of $f$ is the same over $SO(3)$ as over $\operatorname{conv}(SO(3))$. ∎
4.3 A stronger convex relaxation
The results of the previous section show that we need to incorporate constraints allowing us to find solutions in $SO(3)$. The direct incorporation of determinant constraints does not lead to a QCQP, and it is unclear how to compute duals. However, similarly to the previous section we can “solve” single term problems over $SO(3)$ (or at least in the respective convex hull). We therefore introduce auxiliary variables $Q_{ij}=R_{i}R_{j}^{\top}$ and consider the problem
$\min_{\mathbf{R},\mathbf{Q}}$ $-\langle\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle$ (16)
s.t $R_{i}R_{i}^{\top}={\mathtt{I}},$ (19)
$Q_{ij}=R_{i}R_{j}^{\top},$
$Q_{ij}\in SO(3).$
Next we introduce dual variables $\Upsilon_{ii}$ and $\Upsilon_{ij}$ for the constraints (19) and (19) respectively, but we retain (19) as an explicit constraint. The matrices $\Upsilon_{ii}$ are symmetric, $\Upsilon_{ij}=\Upsilon_{ji}^{\top}$, and we view them as blocks of a symmetric matrix $\mathbf{\Upsilon}$. This yields the Lagrangian
$L$ $(\mathbf{R},\mathbf{Q},\mathbf{\Upsilon})=-\langle\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle+\sum\nolimits_{i}\langle\Upsilon_{ii},R_{i}R_{i}^{\top}-{\mathtt{I}}\rangle$
$\qquad\qquad\quad+\sum\nolimits_{i\neq j}\langle\Upsilon_{ij},R_{i}R_{j}^{\top}-Q_{ij}\rangle$
$=\langle\mathbf{\Upsilon}-\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle-\operatorname{tr}\left(\mathbf{\Upsilon}\right)-\sum\nolimits_{i\neq j}\langle\Upsilon_{ij},Q_{ij}\rangle.$ (20)
The dual variables decouple the constraints and makes the Lagrangian quadratic with respect to $\mathbf{R}$ and linear (and separable) with respect to $Q_{ij}$. Therefore the dual function
$d(\mathbf{\Upsilon})=\min_{\mathbf{Q}\in SO(3)^{n\times n},\mathbf{R}}L(\mathbf{R},\mathbf{Q},\mathbf{\Upsilon}),$ (21)
can be computed while explicitly forcing $Q_{ij}\in SO(3)$. The minimum value of $\langle\mathbf{\Upsilon}-\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle$ is 00 if $\mathbf{\Upsilon}-\mathbf{N}\succeq 0$ and $\infty$- ∞ otherwise. Let $\mathcal{I}_{SO(3)}$ be the indicator function,
$\mathcal{I}_{SO(3)}(Q_{ij})=\begin{cases}0&Q_{ij}\in SO(3)\\ \infty&Q_{ij}\notin SO(3)\end{cases},$ (22)
then we are able to write
$\begin{split}\min_{Q_{ij}\in SO(3)}\!\!\!\!&-\langle\Upsilon_{ij},Q_{ij}\rangle\\ &=-\!\left(\max\nolimits_{Q_{ij}}\langle\Upsilon_{ij},Q_{ij}\rangle-\mathcal{I}_{SO(3)}(Q_{ij})\right).\end{split}$ (23)
The last term can be identified as $-\mathcal{I}^{*}_{SO(3)}(\Upsilon_{ij})$ where $\mathcal{I}^{*}_{SO(3)}$ is the convex conjugate [34] of $\mathcal{I}_{SO(3)}$. We thus obtain the dual problem
$\begin{split}\max_{\mathbf{\Upsilon}}&-\operatorname{tr}(\mathbf{\Upsilon})-\sum\nolimits_{i\neq j}\mathcal{I}_{SO(3)}^{*}(\Upsilon_{ij})\\ \text{s.t. }&\mathbf{\Upsilon}-\mathbf{N}\succeq 0.\end{split}$ (24)
Taking the dual once more we get
$\min_{\mathbf{X}\succeq 0}\max_{\mathbf{\Upsilon}}-\operatorname{tr}(\mathbf{\Upsilon})+\langle\mathbf{X},\mathbf{\Upsilon}-\mathbf{N}\rangle-\sum_{i\neq j}\mathcal{I}_{SO(3)}^{*}(\Upsilon_{ij}),$ (25)
which is the same as
$\begin{split}\min_{\mathbf{X}\succeq 0}\max_{\mathbf{\Upsilon}}-\operatorname{tr}(\mathbf{N}\mathbf{X})+\sum\nolimits_{i}\max_{\Upsilon_{ii}}\langle(X_{ii}-{\mathtt{I}}),\Upsilon_{ii}\rangle\\ +\sum\nolimits_{i\neq j}\max_{\Upsilon_{ij}}\left(\langle X_{ij},\Upsilon_{ij}\rangle-\mathcal{I}_{SO(3)}^{*}(\Upsilon_{ij})\right).\end{split}$ (26)
If $X_{ii}\neq{\mathtt{I}}$ it is clear that the maximum over $\Upsilon_{ii}$ will be unbounded. The second term is $\mathcal{I}_{SO(3)}^{**}(X_{ij})$, which is the convex envelope of the indicator function $\mathcal{I}_{SO(3)}$. This is also the indicator function of $\operatorname{conv}(SO(3))$ [34]. Therefore the bidual program—and consequently our proposed relaxation—is given by
$\begin{split}\min_{\mathbf{X}\succeq 0}&-\operatorname{tr}(\mathbf{N}\mathbf{X})\\ \text{s.t. }&X_{ii}={\mathtt{I}},\;X_{ij}\in\operatorname{conv}(SO(3)).\end{split}$ (SDP-cSO(3))
The constraint $X_{ij}\in\operatorname{conv}(SO(3))$ has been shown to be equivalent to a semidefinite constraint [38, 37]. A $3\times 3$ matrix $Y$ is in $\operatorname{conv}(SO(3))$ if and only if
$\mathcal{A}(Y)+{\mathtt{I}}\succeq 0,$ (27)
where $\mathcal{A}(Y)=$
{+pmatrix}$[colsep=0pt,rowsep=0.5pt]-Y_{11}-Y_{22}+Y_{33}&Y_{13}+Y_{31}Y_{12}-Y_{21}Y_{23}+Y_{32}\\ Y_{13}+Y_{31}Y_{11}-Y_{22}-Y_{33}Y_{23}-Y_{32}Y_{12}+Y_{21}\\ Y_{12}-Y_{21}Y_{23}-Y_{32}Y_{11}+Y_{22}+Y_{33}Y_{31}-Y_{13}\\ Y_{23}+Y_{32}Y_{12}+Y_{21}Y_{31}-Y_{13}-Y_{11}+Y_{22}-Y_{33}.$
In contrast to the relaxation in (SDP-O(3)-iso), the inclusion of the convex hull constraints can rule out solutions with incorrect determinants, as we have shown previously. This is crucial when having indefinite cost matrices. While a theoretical guarantee of a tight relaxation is beyond the scope of this paper, it is clear by construction that our new relaxation will be stronger than (SDP-O(3)-iso). In addition, our empirical results show that there is a significant difference between the two approaches. Figure 2 shows the result of applying our new convex relaxation to the synthetic problem described in Section 4.2. In all problem instances the new relaxation returns a solution that is of rank $3$ and therefore optimal in the primal problem.
5 Experiments
We implemented the SDP program using the conic splitting solver [31] with the JuMP [24] wrapper in Julia [^4] Code can be found at: https://ylochman.github.io/anisotropic-ra. The cost matrix in (SDP-cSO(3)) is down-scaled by the average (across the observed relative poses in the scene) of the largest eigenvalue of the computed Hessian matrix, which proved beneficial empirically. The absolute and relative feasibilities are set to $10^{-5}$ and $10^{-6}$, respectively, and infeasibility tolerance is $10^{-8}$. The number of iterations is limited to 500 000500000500\,000500 000. In practice, the methods converge in far fewer iterations to meet the stopping criterion. While it is likely that dedicated solvers are much more efficient, we remark that generalization of efficient methods such as [14] may not be straightforward since they rely heavily on properties $O(d)$, which are not sufficient to ensure $SO(3)$ membership.
(a) $p=0.2$
(b) $p=0.4$
(c) $p=0.6$
(d) $p=0.8$
Figure 3: Histograms of rotation errors wrt. ground truth (top). Corresponding solver runtimes (s) wrt. number of cameras, for different fractions $p$ of observed relative rotations (bottom). Scatter plots represent all instances, and the solid lines are the respective medians.
We compare [^5] Complementing results on both synthetic and real data can be found in the supplementary material. three approaches: 1) SDP-O(3)-iso is the standard rotation averaging approach that uses the regular chordal distance and ignores the $\operatorname{conv}(SO(3))$ constraints, 2) SDP-O(3)-aniso uses the anisotropic objective function but ignores the $\operatorname{conv}(SO(3))$ constraints, and 3) SDP-cSO(3) is the proposed method that uses both the anisotropic objective and the $\operatorname{conv}(SO(3))$ constraints. For completeness of the synthetic study, we also include SDP-cSO(3)-iso, which uses the isotropic objective and the $\operatorname{conv}(SO(3))$ constraints. The reported rotation error wrt. ground truth is the difference $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$, where $R_{i}$ are the ground truth absolute rotations and $R_{i}^{*}$ are the estimated ones. Note that to remove gauge freedom, we align the two sets of rotations $\{R_{i}\}_{i}$ and $\{R_{i}^{*}\}_{i}$, by applying a global rotation $V$ to all rotations in $\{R_{i}\}_{i}$ such that the rotation error achieves its minimum wrt. $V$.
5.1 Synthetic experiments
To create synthetic data, we randomly generate absolute rotations $R_{i}$ and (invertible) covariance matrices $H_{ij}^{-1}$ with eigenvalues in the range $[0.01,0.1]$. For each instance of the problem, we then draw $\Delta\bm{\omega}_{ij}$ from $\mathcal{N}(0,H_{ij}^{-1})$ and compute noisy relative rotations $\tilde{R}_{ij}=e^{[\Delta\bm{\omega}_{ij}]_{\times}}R_{i}R_{j}^{\top}$.
In Figure 3 we evaluate the effect of using the proposed anisotropic error measurements versus regular chordal distances together with the proposed relaxation versus the standard one. We vary the number of cameras from $2$ to $100$, and the proportion of missing data from $0\%$ to $90\%$. For each configuration pair, we run $100$ instances of the problem. In the top row of Figure 3, we report the distance $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$. As expected, using the proposed method (i.e., both new objective and new relaxation) consistently leads to lower errors. We record the time-to-solution of the solver, shown in the bottom row of Figure 3. In this synthetic setting, the proposed relaxation is more efficient. Our intuition is that the introduction of local constraints—$X_{ij}\in\text{conv}(SO(3))$—reduces the oscillations in the employed first-order SDP solver and therefore leads to faster time-to-solution for optimizing both isotropic and anisotropic costs with the proposed constraints.
5.2 Real experiments
[…]
We report the obtained rank for each method and the distance $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$, where the ground truth $\{R_{i}\}_{i}$ is from the final reconstruction obtained using the pipeline of [29]. When SDP-O(3)-aniso does not give a solution $X^{*}$ of rank 3, we compute the closest rank 3 approximation to $X^{*}$ by setting all but the three largest eigenvalues to 00. We then compute a factorization $X^{*}=VV^{\top}$, where $V$ has $3$ columns. From each consecutive $3\times 3$ block $V_{i}$ of $V$ we then compute the closest rotation matrix which we take to be the approximate solution $R_{i}^{*}$.
The proposed method SDP-cSO(3) gives more accurate estimations in all but one of the datasets. We remark that there is no guarantee that anisotropy (or a more correct statistical model in general) will result in better estimates for every noise realization, and all we can expect from MLE is that—on average—it will yield better results. We only have access to a single noise realization for real datasets, and the “ground-truth” solution generated by SfM software may be biased. Our approach is also as fast as the standard method SDP-O(3)-iso in half of the datasets. The results also confirm that the standard semidefinite relaxation fails on real data when anisotropic costs are used (SDP-O(3)-aniso).
Dataset Method $\operatorname{rank}(X^{*})$ $\sqrt{\sum_{i}\|R_{i}\!-\!R_{i}^{*}\|_{F}^{2}}$ Runtime, s
LU Sphinx (70 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.0944 2
SDP-O(3)-aniso 7 18.6037 460
SDP-cSO(3) 3 0.0740 5
Round Church (92 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1399 6
SDP-O(3)-aniso 6 26.3808 632
SDP-cSO(3) 3 0.1267 55
UWO (114 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.3142 14
SDP-O(3)-aniso 6 22.6873 1929
SDP-cSO(3) 3 0.2274 7
Tsar Nikolai I (89 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1170 7
SDP-O(3)-aniso 6 26.8944 1245
SDP-cSO(3) 3 0.0534 5
Vercingetorix (69 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.3146 2
SDP-O(3)-aniso 6 14.8244 242
SDP-cSO(3) 3 0.2910 4
Eglise Du Dome (69 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.0546 4
SDP-O(3)-aniso 6 20.1954 1840
SDP-cSO(3) 3 0.0487 5
King’s College (77 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1656 4
SDP-O(3)-aniso 6 17.6508 681
SDP-cSO(3) 3 0.0796 83
Kronan (131 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.2149 18
SDP-O(3)-aniso 6 22.6035 2997
SDP-cSO(3) 3 0.3892 201
Alcatraz (133 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1763 19
SDP-O(3)-aniso 6 22.7623 1931
SDP-cSO(3) 3 0.1268 107
Museum Barcelona (133 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.2255 22
SDP-O(3)-aniso 7 30.3428 669
SDP-cSO(3) 3 0.1310 27
Temple Singapore (157 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.2646 31
SDP-O(3)-aniso 6 26.0824 2313
SDP-cSO(3) 3 0.1696 628
Table 1: Results on real data from [15, 29]. SDP-O(3)-iso uses the regular chordal distance and ignores the $\operatorname{conv}(SO(3))$ constraints, SDP-O(3)-aniso uses the proposed anisotropic objective but ignores the $\operatorname{conv}(SO(3))$ constraints, SDP-cSO(3) is the proposed approach that uses both the anisotropic objective and the $\operatorname{conv}(SO(3))$ constraints.
6 Conclusions
In this work, we propose a way to incorporate anisotropic costs in a certifiably optimal rotation averaging framework. We demonstrate how traditional solvers designed to optimize chordal distances fail to optimize the new objective due to ignoring the determinant constraint. Our new SDP formulation enforces $\operatorname{conv}(SO(3))$ and our empirical evaluation shows that this relaxation is able to recover global minima. We evaluated the new approach on both synthetic and real datasets and obtained more accurate reconstructions than isotropic methods in most cases, confirming recent observations based on local optimization. We believe that this work will provide a good entry for research into anisotropic rotation averaging, a subarea of rotation averaging that has mostly been set aside because of the focus on isotropic cases thus far.
This work largely focuses on the modeling side and understanding the shortcomings of existing relaxations. We aim to leverage standard SDP solvers; however, the increased strength of our SDP formulation often comes at the cost of more expensive optimization. Designing a dedicated efficient algorithm is an important future direction.
Supplementary Material
Appendix A First-Order Uncertainty Propagation
We leverage Laplace’s approximation, which allows us to identify the Hessian of a non-linear least-squares minimization task as the precision matrix of the posterior $p(\tilde{\bm{\omega}}_{ij}|\text{image $i$ and $j$})$ (up to a positive scale). In the following we apply first-order uncertainty propagation to move from the minimal axis-angle parametrization $\bm{\omega}\in\mathbb{R}^{3}$ to a direct parametrization of rotation matrices $R_{ij}\in\mathbb{R}^{3\times 3}$.
For $\Delta\bm{\omega}\approx\mathbf{0}$ we have $\exp([\Delta\bm{\omega}]_{\times})\approx{\mathtt{I}}+[\Delta\bm{\omega}]_{\times}$. The Jacobian of the map $\Delta\bm{\omega}\to\operatorname{vec}({\mathtt{I}}+[\Delta\bm{\omega}]_{\times})$ is the 9x3 matrix
$\mathtt{J}=\left(\begin{smallmatrix}0&0&0\\ 0&0&1\\ 0&-1&0\\ 0&0&-1\\ 0&0&0\\ 1&0&0\\ 0&1&0\\ -1&0&0\\ 0&0&0\end{smallmatrix}\right).$
Therefore the Gaussian transformed by the mapping $\Delta\bm{\omega}\to\operatorname{vec}({\mathtt{I}}+[\Delta\bm{\omega}]_{\times})$ has covariance matrix
$\mathtt{M}^{-1}=\Sigma_{[\Delta\bm{\omega}]_{\times}}=\mathtt{J}\Sigma_{\Delta\bm{\omega}}\mathtt{J}^{\top}=\mathtt{J}\mathtt{H}^{-1}\mathtt{J}^{\top}.$ (A.1)
Since $\mathtt{J}$ has rank 3, the above relation only fixes 6 out of the 45 degrees of freedom in $\Sigma_{[\Delta\bm{\omega}]_{\times}}$, hence there are 39 d.o.f. available in $\Sigma_{[\Delta\bm{\omega}]_{\times}}$ to achieve the following properties: (i) $\Sigma_{[\Delta\bm{\omega}]_{\times}}$ is invertible and (ii) the maximum-likelihood objective
$\begin{split}&\operatorname{vec}(R-\tilde{R})^{\top}\mathtt{M}\operatorname{vec}(R-\tilde{R})\\ &\doteq\operatorname{vec}(R)^{\top}\mathtt{M}\operatorname{vec}(R)-2\operatorname{vec}(\tilde{R})^{\top}\mathtt{M}\operatorname{vec}(R)\end{split}$ (A.2)
is linear in $R$ and matches $\Delta\bm{\omega}^{\top}\mathtt{H}\Delta\bm{\omega}$ to first order. We consider
$\begin{split}\Delta\bm{\omega}^{\top}\mathtt{H}\Delta\bm{\omega}&\stackrel{{\scriptstyle!}}{{=}}\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}\mathtt{M}[\Delta\bm{\omega}]_{\times})\\ {}&=\operatorname{vec}([\Delta\bm{\omega}]_{\times})^{\top}\operatorname{vec}(\mathtt{M}[\Delta\bm{\omega}]_{\times})\\ {}&=(\mathtt{J}\Delta\bm{\omega})^{\top}\operatorname{vec}(\mathtt{M}[\Delta\bm{\omega}]_{\times})\\ {}&=\Delta\bm{\omega}^{\top}\mathtt{J}^{\top}({\mathtt{I}}\otimes\mathtt{M})\operatorname{vec}([\Delta\bm{\omega}]_{\times})\\ {}&=\Delta\bm{\omega}^{\top}\mathtt{J}^{\top}({\mathtt{I}}\otimes\mathtt{M})\mathtt{J}\Delta\bm{\omega}\end{split}$ (A.3)
for all $\Delta\bm{\omega}$, which implies
$\mathtt{J}^{\top}({\mathtt{I}}\otimes\mathtt{M})\mathtt{J}=\mathtt{J}^{\top}\begin{pmatrix}\mathtt{M}&&\\ &\mathtt{M}&\\ &&\mathtt{M}\end{pmatrix}\mathtt{J}=\mathtt{H}.$ (A.4)
By e.g. using a CAS the relations between $\mathtt{M}$ and $\mathtt{H}$ can be derived as
$\mathtt{M}_{11}+\mathtt{M}_{22}=\mathtt{H}_{33}$ $\mathtt{M}_{11}+\mathtt{M}_{33}=\mathtt{H}_{22}$ $\mathtt{M}_{22}+\mathtt{M}_{33}=\mathtt{H}_{11}$ (A.5)
and $\mathtt{M}_{ij}=-\mathtt{H}_{ij}$ for $i\neq j$. These relations can be more compactly written as
$\mathtt{H}=\operatorname{tr}(\mathtt{M}){\mathtt{I}}-\mathtt{H}$ (A.6)
in accordance with Section 4.1. We relate $[\Delta\bm{\omega}]_{\times}$ and $R$ via $[\Delta\bm{\omega}]_{\times}\approx R\tilde{R}^{\top}-{\mathtt{I}}$ and therefore obtain
$\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}$ $\mathtt{M}[\Delta\bm{\omega}]_{\times})\approx\operatorname{tr}\!\left((R\tilde{R}^{\top}-{\mathtt{I}})^{\top}\mathtt{M}(R\tilde{R}^{\top}-{\mathtt{I}})\right)$
$=\operatorname{tr}(\tilde{R}R^{\top}\mathtt{M}R\tilde{R}^{\top})-2\operatorname{tr}(\mathtt{M}R\tilde{R}^{\top})+\operatorname{tr}(\mathtt{M})$
$=\operatorname{tr}(\mathtt{M}R\tilde{R}^{\top}\tilde{R}R^{\top})-2\operatorname{tr}(\tilde{R}^{\top}\mathtt{M}R)+\operatorname{tr}(\mathtt{M})$
$=2\operatorname{tr}(\mathtt{M})-2\operatorname{tr}(\tilde{R}^{\top}\mathtt{M}R)$
$=2\operatorname{tr}(\mathtt{M})-2\left\langle\mathtt{M}\tilde{R},R\right\rangle_{F},$ (A.7)
which is equivalent to the cost (9) in the main text.
A seemingly different solution $\mathtt{M}^{\prime}$ can be obtained by factoring $\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}\mathtt{M}[\Delta\bm{\omega}]_{\times})$ differently,
$\begin{split}\Delta\bm{\omega}^{\top}\mathtt{H}\Delta\bm{\omega}&\stackrel{{\scriptstyle!}}{{=}}\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}\mathtt{M}^{\prime}[\Delta\bm{\omega}]_{\times})\\ {}&=\operatorname{vec}(\mathtt{M}^{\prime}[\Delta\bm{\omega}]_{\times})^{\top}\operatorname{vec}([\Delta\bm{\omega}]_{\times})\\ {}&=\operatorname{vec}(\mathtt{M}^{\prime}[\Delta\bm{\omega}]_{\times})^{\top}\mathtt{J}\Delta\bm{\omega}\\ {}&=\left((\mathtt{M}^{\prime}\otimes{\mathtt{I}})\operatorname{vec}([\Delta\bm{\omega}]_{\times})\right)^{\top}\mathtt{J}\Delta\bm{\omega}\\ {}&=\Delta\bm{\omega}^{\top}\mathtt{J}^{\top}(\mathtt{M}^{\prime}\otimes{\mathtt{I}})\mathtt{J}\Delta\bm{\omega}\end{split}$ (A.8)
for all $\Delta\bm{\omega}$, leading to the condition $\mathtt{J}^{\top}(\mathtt{M}^{\prime}\otimes{\mathtt{I}})\mathtt{J}=\mathtt{H}$. Using a CAS it can be seen that these conditions on $\mathtt{M}^{\prime}$ are the same as for $\mathtt{M}$ in (A.5), and therefore $\mathtt{M}=\mathtt{M}^{\prime}$.
Appendix B Parameterization of $R$
There are different ways of parametrizing rotations. In the main paper we use the $R=e^{[\Delta\bm{\omega}]_{\times}}\tilde{R}$ which lead to the objective of the form $-\langle M\tilde{R},R\rangle$. In this section, we investigate the effects of switching to other possible parameterizations. In particular, we start by letting $R=e^{[\bm{\omega}]_{\times}}=e^{[\tilde{\bm{\omega}}+\Delta\bm{\omega}]_{\times}}$ and look at its approximations.
We first notice that we can expect $R\tilde{R}^{\top}\approx{\mathtt{I}}$ as $\tilde{R}$ is a noisy realization of $R$, therefore $R$ and $\tilde{R}^{\top}$ approximately commute and
$\begin{split}e^{[\Delta\bm{\omega}]_{\times}}&=e^{[\bm{\omega}-\tilde{\bm{\omega}}]_{\times}}=e^{-\alpha[\tilde{\bm{\omega}}]_{\times}+[\bm{\omega}]_{\times}-(1-\alpha)[\tilde{\bm{\omega}}]_{\times}}\\ &\approx e^{-\alpha[\tilde{\bm{\omega}}]_{\times}}e^{[\bm{\omega}]_{\times}}e^{-(1-\alpha)[\tilde{\bm{\omega}}]_{\times}}\\ &=(\tilde{R}^{\top})^{\alpha}R(\tilde{R}^{\top})^{1-\alpha}\end{split}$ (B.9)
for any $\alpha\in[0,1]$, i.e.
$R\approx\tilde{R}^{\alpha}e^{[\Delta\bm{\omega}]_{\times}}\tilde{R}^{(1-\alpha)}.$ (B.10)
The Hessian induced by the two-view optimization can then be computed in accordance with this mapping. The first-order Taylor expansion of (B.9) results in
$[\Delta\bm{\omega}]_{\times}$ $\approx(\tilde{R}^{\top})^{\alpha}R(\tilde{R}^{\top})^{1-\alpha}-{\mathtt{I}}.$ (B.11)
Plugging this into (5) gives the corresponding linear cost
$-\langle\tilde{R}^{\alpha}M\tilde{R}^{(1-\alpha)},R\rangle.$ (B.12)
The natural choices for $\alpha$ are 00, $1/2$ and $1$. Setting $\alpha=0$ yields the formulation employed in the main paper. We found that for the other values of $\alpha$ the solution of anisotropic rotation averaging is the same. However, when using the Hessian matrix computed from the Jacobian of the initial parameterization, $R=e^{[\tilde{\bm{\omega}}+\Delta\bm{\omega}]_{\times}}$, the results vary when optimizing the anisotropic cost for different values of $\alpha$. In our synthetic experiments the best solution is often obtained with $\alpha=1/2$, however, this setting does not outperform the proposed formulation overall.
Appendix C Experimental Details
We provide complementing results on synthetic and real datasets below.
C.1 Synthetic experiments
Figure C.1: The configuration of the cameras (top left). Rotation error wrt. ground truth $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$ (top right, bottom) for the increasing noise scale around one of the axes.
We set up a synthetic graph with three cameras as shown in Figure C.1. The estimated relative rotations corresponding to the black dashed edges are “certain”, i.e. $\Delta\bm{\omega}_{ij}$ are drawn from $\mathcal{N}(0,\varepsilon{\mathtt{I}})$, where $\varepsilon=0.001$. The estimated relative rotation of the gray edge has varying uncertainty around one of the three axes, e.g., for the x- (red) axis, the noise covariance is $\operatorname{diag}(\sigma,\varepsilon,\varepsilon)$, where $\sigma$ varies from $0.01$ to $0.3$. The effects on the error are shown in Figure C.1. The proposed method relies on certain relative rotations and gives an accurate solution, while the standard isotropic approach is negatively affected by the single noisy relative rotation.
(a) $p=0.1$
(b) $p=0.3$
(c) $p=0.5$
(d) $p=0.7$
(e) $p=0.9$
(f) $p=1.0$
Figure C.2: Histograms of rotation errors wrt. ground truth (rows 1,3,5). Corresponding solver runtime (s) wrt. the increasing number of cameras (rows 2,4,6). Results are shown for different fractions $p$ of observed relative rotations.
We also present complementing results of the synthetic experiments of Sec 5.1 in the paper. Figure C.2 shows the rotation error histograms and the dependency of the runtime on the number of cameras for the other fractions of the observed relative rotations. These results are in line with the analysis provided in the paper — using the proposed objective leads to lower errors and using the proposed constraints speeds up the employed SDP solver.
C.2 Real experiments
To account for the estimated uncertainties in the evaluation, we compared the methods using the Mahalanobis distance between the axis-angle vectors of the estimated and ground truth rotations. The axis-angle vector $\bm{\omega}_{i}$ is distributed according to $\mathcal{N}(\bm{\omega}_{i}^{*},H_{i})$, where $\bm{\omega}_{i}^{*}$ is the ground truth axis-angle vector and $H_{i}=\sum_{j:\{ij\}\text{ was observed}}H_{ij}$, assuming that all other cameras are fixed. Let $\Delta\bm{\omega}_{i}^{-}=\bm{\omega}_{i}-\bm{\omega}_{i}^{*}$ and $\Delta\bm{\omega}_{i}^{+}=\bm{\omega}_{i}+\bm{\omega}_{i}^{*}$, then the Mahalanobis error is
$\sqrt{\sum_{i}\min\{\Delta\bm{\omega}_{i}^{-\top}H_{i}\Delta\bm{\omega}_{i}^{-},\Delta\bm{\omega}_{i}^{+\top}H_{i}\Delta\bm{\omega}_{i}^{+}\}},$ (C.13)
where the minimization is done to account for the sign ambiguity in the axis-angle representation. We also present the RMS angular errors for a better geometric interpretation. As shown in Table C.1, in many cases, the proposed method leads to a much lower error.
Dataset Method Mahal. err. Angl. err.
LU Sphinx SDP-O(3)-iso 0.3879 0.46
SDP-cSO(3) 0.2067 0.36
Round Church SDP-O(3)-iso 0.6308 0.59
SDP-cSO(3) 0.3683 0.54
UWO SDP-O(3)-iso 1.4814 1.19
SDP-cSO(3) 0.7267 0.86
Tsar Nikolai I SDP-O(3)-iso 0.6871 0.48
SDP-cSO(3) 0.1878 0.22
Vercingetorix SDP-O(3)-iso 0.4312 1.53
SDP-cSO(3) 0.4228 1.42
Eglise Du Dome SDP-O(3)-iso 0.2239 0.24
SDP-cSO(3) 0.1882 0.21
King’s College SDP-O(3)-iso 0.2288 0.76
SDP-cSO(3) 0.1303 0.37
Kronan SDP-O(3)-iso 0.7380 0.76
SDP-cSO(3) 1.1114 1.38
Alcatraz SDP-O(3)-iso 1.3329 0.62
SDP-cSO(3) 1.0110 0.45
Museum Barcelona SDP-O(3)-iso 2.7101 0.79
SDP-cSO(3) 1.2164 0.46
Temple Singapore SDP-O(3)-iso 2.4197 0.86
SDP-cSO(3) 1.0760 0.55
Table C.1: Mahalanobis distance between the axis-angle vectors of the estimated and ground truth rotations and RMS angular errors (degrees) evaluated on the real datasets.

### v2

@@
 However, global optimization for rotation averaging has remained a challenge in this scenario.
-In this paper we show how anisotropic costs can be incorporated in certifiably optimal rotation averaging.
+In this work we show how anisotropic costs can be incorporated in certifiably optimal rotation averaging.
 We also demonstrate how existing solvers, designed for isotropic situations, fail in the anisotropic setting.
-Finally, we propose a stronger relaxation and show empirically that it is able to recover global optima in all tested datasets and leads to a more accurate reconstruction in all but one of the scenes.
+Finally, we propose a stronger relaxation and empirically show that it recovers global optima in all tested datasets and leads to more accurate reconstructions in almost all scenes.
+[^1] This work was supported by the Wallenberg Artificial Intelligence, Autonomous Systems and Software Program (WASP), funded by the Knut and Alice Wallenberg Foundation, and the Swedish Research Council (#2023-05341).
 1 Introduction
@@
 The key strength here is that rotation averaging optimizes the orientations of all the cameras at the same time, allowing recovery methods to avoid incrementally increasing errors (drift) [14] that easily occur in sequential SfM methods.
-More recently, [5, 7, 8] also point out the importance of rotation averaging for improving accuracy and speed of simultaneous localization and mapping (SLAM).
+More recently, [7, 9, 10] point out the importance of rotation averaging to improve the accuracy and speed of simultaneous localization and mapping (SLAM).
 The main difficulty in solving the rotation averaging problem is that rotations reside in a nonlinear manifold $SO(3)$.
@@
 In this paper we will only consider matrix representations.
-The works [2, 21] also considers certifiable optimization of the anisotropic chordal distance.
+The works [3, 23] also consider certifiable optimization of the anisotropic chordal distance.
 However, their SDP relaxation is based on the Cayley mapping and the relation to our formulation is somewhat unclear.
@@
 It is likely, in the light of our analysis in this paper, and the discussion in [6] that this is a result of the authors only enforcing $O(3)$ membership for their optimization.
-It is also worth mentioning that—while this paper looks at the anisotropic $L_{2}$ chordal distance—, it is possible to consider other distances between rotations as well as their robust counterparts.
-E.g., [43, 44] consider the geodesic distance (i.e.
-the one on the manifold of $SO(3)$).
+It is also worth mentioning that—while this paper looks at the anisotropic $L_{2}$ chordal distance—it is possible to consider other distances between rotations as well as their robust counterparts.
+E.g., [47, 48] consider the geodesic distance.
 Further, in [11, 43, 46] the authors analyze and look at different (robustified) distances in order to obtain a “robust” formulation (resulting in the solution being less susceptible to outliers).
 Outliers are handled via the introduction of auxiliary variables in the SDP formulation in [38, 8].
-From an algorithmic point of view, the SPD relaxation can be solved by general-purpose solvers, e.g.
+From an algorithmic point of view, the SDP relaxation can be solved by general-purpose solvers, e.g.
 [1].
@@
 The authors show that such a method can be made highly effective.
+Like most work on rotation averaging, we focus on point estimates and assume unimodal Gaussian noise in the observations.
+If the input data has multiple modes (e.g.
+due to perceptual aliasing), the particle-based algorithm proposed in [4] provides more expressive marginals.
+Perceptual aliasing can also be explicitly addressed by leveraging cycle consistency constraints (e.g.
+[50, 25, 42]).
 3 Certifiably Optimal Rotation Averaging
@@
 $\min_{R_{i}\in SO(3)}\sum_{i\neq j}\|\tilde{R}_{ij}-R_{i}R_{j}^{\top}\|_{F}^{2}.$ (1)
-The derivation of a strong convex relaxation relies on two observations: Firstly, since the set of rotation matrices have constant (Frobenius) norm, the objective function can be replaced by $-\sum_{i\neq j}\langle\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle$.
-If we let $\mathbf{R}=\begin{bmatrix}R_{1}^{\top}&R_{2}^{\top}&\ldots\end{bmatrix}^{\top}$ and $\mathbf{\tilde{R}}$ be the matrix containing the blocks $\tilde{R}_{ij}$ (and zeros where no relative rotation estimate is available) we can simplify the objective to $-\langle\mathbf{\tilde{R}},\mathbf{R}\mathbf{R}^{\top}\rangle$, which is linear in the relative rotations.
+The derivation of a strong convex relaxation relies on two observations: Firstly, since the set of rotation matrices have constant (Frobenius) norm, the objective function can be replaced by $\sum_{i\neq j}\langle\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle$.
+If we let $\mathbf{R}=\begin{bmatrix}R_{1}^{\top}&R_{2}^{\top}&\ldots\end{bmatrix}^{\top}$ and $\mathbf{\tilde{R}}$ be the matrix containing the blocks $\tilde{R}_{ij}$ (and zeros where no relative rotation estimate is available) we can simplify the objective to $\langle\mathbf{\tilde{R}},\mathbf{R}\mathbf{R}^{\top}\rangle$, which is linear in the relative rotations.
 Secondly, by orthogonality we have $R_{i}R_{i}^{\top}={\mathtt{I}}$, resulting in the Quadratically Constrained Quadratic Program (QCQP)
@@
 QCQP are typically non-convex but can be relaxed to convex linear semidefinite programs (SDP) by replacing $\mathbf{R}\mathbf{R}^{\top}$ with a positive semidefinite matrix $\mathbf{X}$ (technically by taking the Lagrange dual twice), yielding the relaxation
-[…]
+$d_{2}^{*}:=\min_{\mathbf{X}\succeq 0}$ $\,-\langle\mathbf{\tilde{R}},\mathbf{X}\rangle\quad\text{s.t. }\forall i:X_{ii}={\mathtt{I}}.$ (SDP-O(3)-iso)
 The above program is convex and can therefore be solved reliably using general purpose [1, 26] or specialized solvers [16, 5, 18].
@@
 Moreover, if the solution $\mathbf{X}$ to (SDP-O(3)-iso) has $\operatorname{rank}(\mathbf{X})=3$ then it can be factorized into $\mathbf{X}=\mathbf{R}\mathbf{R}^{\top}$, with $\mathbf{R}$ being feasible in (2), which yields $p^{*}=d_{2}^{*}$ and certifies that $\mathbf{R}$ is globally optimal in (2).
-Thus, a solution to the original non-convex QCQP is obtained through optimization via convex SDP program.
+Thus, a solution of the original non-convex QCQP is obtained via a convex SDP program.
 We remark that the use of a linear objective is important to obtain a strong relaxation.
@@
 The left image shows the ground truth solution (obtained after bundle adjustment).
-To generate the graphs to the right, we sampled rotations of the second camera around its x-,y-, and z-axes between $-5$ and $5$ degrees.
+To generate the graphs to the right, we sampled rotations of the second camera around its x-,y-, and z-axes between 5-5 and $5$ degrees.
 For each sampled rotation, we found the best 3D point locations and position of the second camera, and computed the sum-of-squared (calibrated) reprojection errors (solid curves).
 The result shows that rotations around the black/upward axis affects the reprojection errors significantly less than the others, resulting in a direction of larger uncertainty.
-By performing the sensitivity analysis (i.e.
+By performing sensitivity analysis (i.e.
 by looking at the singular values of the Jacobian) one can arrive at the similar conclusion.
@@
 4.1 Incorporating uncertainties
-Consider a relative rotation $\tilde{R}$ that has been optimized to a minimum using a nonlinear least squares approach in a local two-view optimization.
-Using the exponential map we write $R=e^{[\Delta\bm{\omega}]_{\times}}\tilde{R}$ [^1] Other parameterizations are discussed in the supplementary material., where $\Delta\bm{\omega}$ is an axis-angle representation of the deviation from $\tilde{R}$, i.e.
-the axis-angle of $R\tilde{R}^{\top}$.
-Since $\tilde{R}$ is a minimum, gradient terms vanish and the two-view objective can locally be approximated by a quadratic function (ignoring an inconsequential constant)
-$\tfrac{1}{2}\Delta\bm{\omega}^{\top}H\Delta\bm{\omega}.$ (3)
-The positive semidefinite matrix $H=J^{\top}J$ (where $J$ is the Jacobian of the residual function) approximates the growth of the objective function in a region around the minimum $\tilde{R}$.
-We remark that the initial two-view objective function can depend on additional parameters such as camera positions and 3D point locations.
-Here we assume that these “nuisance” parameters have been marginalized out and therefore work with a reduced objective solely in terms of the rotation parameters $\Delta\bm{\omega}$ (using the Schur complement to factor out additional unknowns).
-We assume Laplace’s approximation to be valid at least locally in order to quantify the uncertainty of $\Delta\bm{\omega}$, i.e.
-identify the precision matrix with the Hessian $H$.
-While in general the uncertainty of a $3\times 3$ matrix $R$ is represented by a $9\times 9$ covariance/precision matrix, we aim for a linear cost (in terms of rotation matrices) in order to achieve a strong relaxation, which means that the objective is restricted to an inner product between $R$ and some other (input-dependent) matrix.
-We first find a matrix $M$ such that
-$\begin{split}\Delta\bm{\omega}^{\top}H\Delta\bm{\omega}&=\operatorname{tr}\left([\Delta\bm{\omega}]_{\times}^{\top}M[\Delta\bm{\omega}]_{\times}\right)\\ &=-\operatorname{tr}\left(M[\Delta\bm{\omega}]_{\times}^{2}\right).\end{split}$ (4)
-Using the identity $[\Delta\bm{\omega}]_{\times}^{2}=\Delta\bm{\omega}^{\top}\Delta\bm{\omega}{\mathtt{I}}-\Delta\bm{\omega}\Delta\bm{\omega}^{\top}$, we obtain that $H=\operatorname{tr}(M){\mathtt{I}}-M$.
-Taking the trace on both sides further shows that $\operatorname{tr}(H)=2\operatorname{tr}(M)$, resulting in
-$M=\tfrac{\operatorname{tr}(H)}{2}{\mathtt{I}}-H.$ (5)
+Consider a relative rotation $\tilde{R}_{ij}$ between cameras $i$ and $j$, that has been fully optimized using e.g.
+a Gauss-Newton method for two-view optimization.
+Using the exponential map we write $R_{ij}=e^{[\Delta\bm{\omega}_{ij}]_{\times}}\tilde{R}_{ij}$ [^2] Other parameterizations are discussed in the supplementary material., where $\Delta\bm{\omega}_{ij}$ is an axis-angle representation of the deviation from $\tilde{R}_{ij}$, i.e.
+the axis-angle of $R_{ij}\tilde{R}_{ij}^{\top}$.
+Since $\tilde{R}_{ij}$ is a minimum, gradient terms vanish and the two-view objective can locally be approximated by a quadratic function (up to an irrelevant constant)
+$\tfrac{1}{2}\Delta\bm{\omega}_{ij}^{\top}H_{ij}\Delta\bm{\omega}_{ij}.$ (3)
+We assume that $H_{ij}$ is positive semidefinite—usually given by a Gauss-Newton approximation $H_{ij}=J_{ij}^{\top}J_{ij}$, where $J_{ij}$ is the Jacobian of the underlying residual function.
+We remark that the two-view objective may depend on additional parameters such as camera positions and 3D point locations.
+Here we assume that these “nuisance” parameters have been marginalized out (using the Schur complement), and therefore we work with a reduced objective solely in terms of the rotation parameters $\Delta\bm{\omega}_{ij}$.
+We assume Laplace’s approximation to be valid at least locally in order to quantify the uncertainty of $\Delta\bm{\omega}_{ij}$, i.e.
+identify the precision matrix with the Hessian $H_{ij}$.
+While in general the uncertainty of a $3\times 3$ matrix $R_{ij}$ is represented by a $9\times 9$ covariance/precision matrix, we aim for a linear cost (in terms of rotation matrices) in order to achieve a strong relaxation, which means that the objective is restricted to an inner product between $R_{ij}$ and a constant but input-dependent matrix.
+We first determine a matrix $M_{ij}$ such that
+$\begin{split}\Delta\bm{\omega}_{ij}^{\top}H_{ij}\Delta\bm{\omega}_{ij}&=\operatorname{tr}\left([\Delta\bm{\omega}_{ij}]_{\times}^{\top}M_{ij}[\Delta\bm{\omega}_{ij}]_{\times}\right)\\ &=-\operatorname{tr}\left(M_{ij}[\Delta\bm{\omega}_{ij}]_{\times}^{2}\right).\end{split}$ (4)
+Using the identity $\mathbf{v}]_{\times}^{2}=\mathbf{v}^{\top}\mathbf{v}{\mathtt{I}}-\mathbf{v}\mathbf{v}^{\top}$, we obtain that $H_{ij}=\operatorname{tr}(M_{ij}){\mathtt{I}}-M_{ij}$.
+Taking the trace on both sides further shows that $\operatorname{tr}(H_{ij})=2\operatorname{tr}(M_{ij})$, resulting in
+$M_{ij}=\tfrac{\operatorname{tr}(H_{ij})}{2}{\mathtt{I}}-H_{ij}.$ (5)
 Next, we deduce via the Taylor expansion of the exponential map that
-$[\Delta\bm{\omega}]_{\times}\approx R\tilde{R}^{\top}-{\mathtt{I}},$ (6)
+$\Delta\bm{\omega}_{ij}]_{\times}\approx R_{ij}\tilde{R}_{ij}^{\top}-{\mathtt{I}},$ (6)
 with equality up to first order terms. This yields the objective corresponding to (3) based on (4),
-$\tfrac{1}{2}\operatorname{tr}\big{(}(R\tilde{R}^{\top}-{\mathtt{I}})^{\top}M(R\tilde{R}^{\top}-{\mathtt{I}})\big{)}.$ (7)
+$\tfrac{1}{2}\operatorname{tr}\big{(}(R_{ij}\tilde{R}_{ij}^{\top}-{\mathtt{I}})^{\top}M_{ij}(R_{ij}\tilde{R}_{ij}^{\top}-{\mathtt{I}})\big{)}.$ (7)
 Similarly to the isotropic case, by leveraging the properties of rotation matrices and by omitting constants, this objective can be reduced to a linear one [^3] In the supplementary material we show that we arrive at the same objective using classical first-order uncertainty propagation.,
-$-\operatorname{tr}\big{(}M\tilde{R}R^{\top}\big{)}=-\langle M\tilde{R},R\rangle.$ (8)
-We observe that this objective can be viewed as a negative log-likelihood of a Langevin distribution $\mathcal{L}(R,M)$, e.g. [11], with a density function of the type
-$p(\tilde{R})\propto e^{\langle M\tilde{R},R\rangle},$ (9)
-and that generating rotations $\tilde{R}$ according to (9) is approximately (up to second order terms) equivalent to generating axis-angle vectors $\Delta\bm{\omega}$ from a normal distribution $\mathcal{N}(0,H^{-1})$ with density
-$p(\Delta\bm{\omega})\propto e^{-\frac{1}{2}\Delta\bm{\omega}^{\top}H\Delta\bm{\omega}}.$ (10)
-Hence, the above derivations suggest a way of generalizing the rotation averaging approach by replacing the relative estimates $\tilde{R}_{ij}$ with a weighted version $M_{ij}\tilde{R}_{ij}$, where $M_{ij}$ is computed from the Hessian $H_{ij}$ of the relative pose problem as outlined above.
+$\operatorname{tr}\big{(}M_{ij}\tilde{R}_{ij}R_{ij}^{\top}\big{)}=-\langle M_{ij}\tilde{R}_{ij},R_{ij}\rangle.$ (8)
+We observe that this objective can be viewed as a negative log-likelihood of a Langevin distribution $\mathcal{L}(R_{ij},M_{ij})$, e.g. [13], with a density function of the type
+$p(\tilde{R}_{ij})\propto e^{\langle M_{ij}\tilde{R}_{ij},R_{ij}\rangle},$ (9)
+and that generating rotations $\tilde{R}_{ij}$ according to (9) is approximately (up to second order terms) equivalent to generating axis-angle vectors $\Delta\bm{\omega}_{ij}$ from a normal distribution $\mathcal{N}(0,H^{-1}_{ij})$ with density
+$p(\Delta\bm{\omega}_{ij})\propto e^{-\frac{1}{2}\Delta\bm{\omega}_{ij}^{\top}H_{ij}\Delta\bm{\omega}_{ij}}.$ (10)
+Hence, the above derivations suggest to generalize the rotation averaging approach by replacing the relative estimates $\tilde{R}_{ij}$ with a weighted version $M_{ij}\tilde{R}_{ij}$, where $M_{ij}$ is based on the Hessian $H_{ij}$ of the relative pose problem as outlined above.
 Therefore, our proposed objective to minimize is
-$-\sum\nolimits_{i\neq j}\langle M_{ij}\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle=-\langle\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle,$ (11)
-where $\mathbf{N}$ is a symmetric block matrix containing the blocks $(M_{ij}\tilde{R}_{ij})^{\top}$ for $i<j$, 00 when $i=j$, and $M_{ji}\tilde{R}_{ji}$ for $i>j$.
-We conclude this section by remarking that—in contrast to $H$—the matrix $M$ will generally not be positive semidefinite. The following lemma characterizes its eigenvalues
+$\sum\nolimits_{i\neq j}\langle M_{ij}\tilde{R}_{ij},R_{j}R_{i}^{\top}\rangle=-\langle\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle,$ (11)
+where $\mathbf{N}$ is a symmetric block matrix containing the blocks $(M_{ij}\tilde{R}_{ij})^{\top}$ for $i<j$, 0 when $i=j$, and $M_{ji}\tilde{R}_{ji}$ for $i>j$.
+We conclude this section by remarking that—in contrast to $H_{ij}$—the matrix $M_{ij}$ will generally not be positive semidefinite as the following lemma shows.
 Lemma 1.
-If $H\succeq 0$ then $M=\frac{\operatorname{tr}(H)}{2}{\mathtt{I}}-H$ has eigenvalues $\lambda_{1}\geq\lambda_{2}\geq|\lambda_{3}|$.
+If $H_{ij}\succeq 0$ then $M_{ij}=\frac{\operatorname{tr}(H_{ij})}{2}{\mathtt{I}}-H_{ij}$ has eigenvalues $\lambda_{1}\geq\lambda_{2}\geq|\lambda_{3}|$.
 Proof.
-If $\eta_{1}\geq\eta_{2}\geq\eta_{3}\geq 0$ are eigenvalues of $H$ then $\frac{\operatorname{tr}(H)}{2}-\eta_{i}$, $i=1,2,3$ are the eigenvalues of $M$.
+If $\eta_{1}\geq\eta_{2}\geq\eta_{3}\geq 0$ are eigenvalues of $H_{ij}$ then $\frac{\operatorname{tr}(H_{ij})}{2}-\eta_{i}$, $i=1,2,3$ are the eigenvalues of $M_{ij}$.
 Sorted in decreasing order we get $\lambda_{1}=\frac{1}{2}(\eta_{1}+\eta_{2}-\eta_{3})$, $\lambda_{2}=\frac{1}{2}(\eta_{1}-\eta_{2}+\eta_{3})$ and $\lambda_{3}=\frac{1}{2}(-\eta_{1}+\eta_{2}+\eta_{3})$.
@@
 ∎
-The case $\lambda_{3}<0$ occurs when the leading eigenvalue of $H$ is significantly larger than the other two, i.e.
-the estimate $\bm{\omega}$ is more certain in a specific direction.
+The case $\lambda_{3}<0$ occurs when the leading eigenvalue of $H_{ij}$ is significantly larger than the other two, i.e.
+the estimate $\bm{\omega}_{ij}$ is more certain in a specific direction.
 This is actually a rather common scenario in image-based estimation as the in-plane part of the rotation is usually better constrained than the out-of-plane rotation angles.
+Table 1 shows the percentage of indefinite matrices $M_{ij}$ for the real datasets tested in Section 5.2.
 4.2 Global solutions: $O(3)$ vs. $SO(3)$
@@
 Fig.
-2 shows that the standard rotation averaging relaxation is never able to recover a rank-$3$ solution [^3] To determine rank, we used the smallest number $N$ of singular values (i.e., the first $N$ largest values) that sum up to $>99.9\%$ of the total sum.
+2 shows that the standard rotation averaging relaxation is never able to recover a rank-$3$ solution [^4] To determine rank, we used the smallest number $N$ of singular values (i.e., the first $N$ largest values) that sum up to $99.9\%$ of the total sum.
 for any of the problem instances.
@@
 Figure 2: The results of running 1000 synthetic instances of rotation averaging with an anisotropic objective.
-In contrast to the proposed approach the standard convex relaxations give no solutions of rank $3$.
-To better understand the problem, we consider terms of type $-\langle M\tilde{R},R\rangle$ that constitute the objective function.
+In contrast to the proposed approach the standard convex relaxations give no solutions of rank 33.
+To better understand the problem, we consider terms of type $\langle M\tilde{R},R\rangle$ [^5] For the remainder of this section we focus on a single relative rotation and therefore omit the subscripts.
+that constitute the objective function.
 As we have seen in Section 4.1, these are locally accurate and allow uncertainties of $\tilde{R}$ to be propagated to the rotation averaging stage.
@@
 Proof.
-We minimize $-\langle M\tilde{R},R\rangle$ by computing KKT-points over $O(3)$.
-Introducing a (symmetric) Lagrange multiplier $\Lambda$ for the constraint $RR^{\top}-{\mathtt{I}}=0$ and differentiating
+We minimize $\langle M\tilde{R},R\rangle$ by computing KKT-points over $O(3)$. Introducing a (symmetric) Lagrange multiplier $\Lambda$ for the constraint $RR^{\top}-{\mathtt{I}}=0$ and differentiating
 $L(R,\Lambda)=-\langle M\tilde{R},R\rangle+\tfrac{1}{2}\langle\Lambda,RR^{\top}-{\mathtt{I}}\rangle$ (13)
@@
 with the corresponding objective value
-$-\langle UDU^{\top}\tilde{R},USU^{\top}\tilde{R}\rangle=\pm\lambda_{1}\pm\lambda_{2}\pm\lambda_{3},$ (15)
+$\langle UDU^{\top}\tilde{R},USU^{\top}\tilde{R}\rangle=\pm\lambda_{1}\pm\lambda_{2}\pm\lambda_{3},$ (15)
 where $\lambda_{1},\lambda_{2},\lambda_{3}$ are the eigenvalues of $M$.
 According to Lemma 1 the smallest eigenvalue $\lambda_{3}$ of $M$ is negative if $M$ is indefinite.
-In this case the objective function is minimized over $O(3)$ by choosing $S=\operatorname{diag}(1,1,-1)$ which yields the solution $R=USU^{\top}\tilde{R}$ and $-\langle M\tilde{R},R\rangle=-\lambda_{1}-\lambda_{2}-|\lambda_{3}|$.
-In contrast $-\langle M\tilde{R},\tilde{R}\rangle=-\lambda_{1}-\lambda_{2}+|\lambda_{3}|$, showing that the smallest value of $f$ is $f(R)=-2|\lambda_{3}|$.
-Since $SO(3)\subset O(3)$ the KKT points of $SO(3)$ are those which satisfy $\det(R)=1$, which implies that the matrix $S$ has to have either zero or two elements that are $-1$.
+In this case the objective function is minimized over $O(3)$ by choosing $S=\operatorname{diag}(1,1,-1)$ which yields the solution $R=USU^{\top}\tilde{R}$ and $\langle M\tilde{R},R\rangle=-\lambda_{1}-\lambda_{2}-|\lambda_{3}|$.
+In contrast $\langle M\tilde{R},\tilde{R}\rangle=-\lambda_{1}-\lambda_{2}+|\lambda_{3}|$, showing that the smallest value of $f$ is $f(R)=-2|\lambda_{3}|$.
+Since $SO(3)\subset O(3)$ the KKT points of $SO(3)$ are those which satisfy $\det(R)=1$, which implies that the matrix $S$ has to have either zero or two elements that are 1-1.
 Therefore the above choice of $S$ is infeasible, and the optimal $S$ is instead given by $S=I$, which shows that the minimum value of $f$ is $f(\tilde{R})=0$.
 ∎
-The above result shows that loss-terms of the type $-\langle M\tilde{R},R\rangle$ will favor incorrect solutions (that can be far from the estimation $\tilde{R}$) when optimizing over $O(3)$, which explains the results in Figure 2.
+The above result shows that loss-terms of the type $\langle M\tilde{R},R\rangle$ will favor incorrect solutions (that can be far from the estimation $\tilde{R}$) when optimizing over $O(3)$, which explains the results in Figure 2.
 While strictly enforcing $R\in SO(3)$ is undesirable, as it results in non-convex constraints, the following result shows that it is enough to use $R\in\operatorname{conv}(SO(3))$ to resolve this issue.
@@
 For any set $C$, the extreme points of $\operatorname{conv}(C)$ are always contained in $C$ [37].
-Further, since for every point $Q\in SO(3)$ the linear function $-\langle Q,R\rangle$ has its unique minimum in $R=Q$, it can be deduced that the set of extreme points of $\operatorname{conv}(SO(3))$ is exactly $SO(3)$.
+Further, since for every point $Q\in SO(3)$ the linear function $\langle Q,R\rangle$ has its unique minimum in $R=Q$, it can be deduced that the set of extreme points of $\operatorname{conv}(SO(3))$ is exactly $SO(3)$.
 Thus, the minimum value of $f$ is the same over $SO(3)$ as over $\operatorname{conv}(SO(3))$.
@@
 can be computed while explicitly forcing $Q_{ij}\in SO(3)$.
-The minimum value of $\langle\mathbf{\Upsilon}-\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle$ is 00 if $\mathbf{\Upsilon}-\mathbf{N}\succeq 0$ and $\infty$- ∞ otherwise.
+The minimum value of $\langle\mathbf{\Upsilon}-\mathbf{N},\mathbf{R}\mathbf{R}^{\top}\rangle$ is 0 if $\mathbf{\Upsilon}-\mathbf{N}\succeq 0$ and -$\infty$ otherwise.
 Let $\mathcal{I}_{SO(3)}$ be the indicator function,
@@
 $\begin{split}\min_{Q_{ij}\in SO(3)}\!\!\!\!&-\langle\Upsilon_{ij},Q_{ij}\rangle\\ &=-\!\left(\max\nolimits_{Q_{ij}}\langle\Upsilon_{ij},Q_{ij}\rangle-\mathcal{I}_{SO(3)}(Q_{ij})\right).\end{split}$ (23)
-The last term can be identified as $-\mathcal{I}^{*}_{SO(3)}(\Upsilon_{ij})$ where $\mathcal{I}^{*}_{SO(3)}$ is the convex conjugate [34] of $\mathcal{I}_{SO(3)}$. We thus obtain the dual problem
+The last term can be identified as $\mathcal{I}^{*}_{SO(3)}(\Upsilon_{ij})$ where $\mathcal{I}^{*}_{SO(3)}$ is the convex conjugate [37] of $\mathcal{I}_{SO(3)}$. We thus obtain the dual problem
 $\begin{split}\max_{\mathbf{\Upsilon}}&-\operatorname{tr}(\mathbf{\Upsilon})-\sum\nolimits_{i\neq j}\mathcal{I}_{SO(3)}^{*}(\Upsilon_{ij})\\ \text{s.t.
@@
 $\begin{split}\min_{\mathbf{X}\succeq 0}&-\operatorname{tr}(\mathbf{N}\mathbf{X})\\ \text{s.t. }&X_{ii}={\mathtt{I}},\;X_{ij}\in\operatorname{conv}(SO(3)).\end{split}$ (SDP-cSO(3))
-The constraint $X_{ij}\in\operatorname{conv}(SO(3))$ has been shown to be equivalent to a semidefinite constraint [38, 37]. A $3\times 3$ matrix $Y$ is in $\operatorname{conv}(SO(3))$ if and only if
-$\mathcal{A}(Y)+{\mathtt{I}}\succeq 0,$ (27)
-where $\mathcal{A}(Y)=$
-{+pmatrix}$[colsep=0pt,rowsep=0.5pt]-Y_{11}-Y_{22}+Y_{33}&Y_{13}+Y_{31}Y_{12}-Y_{21}Y_{23}+Y_{32}\\ Y_{13}+Y_{31}Y_{11}-Y_{22}-Y_{33}Y_{23}-Y_{32}Y_{12}+Y_{21}\\ Y_{12}-Y_{21}Y_{23}-Y_{32}Y_{11}+Y_{22}+Y_{33}Y_{31}-Y_{13}\\ Y_{23}+Y_{32}Y_{12}+Y_{21}Y_{31}-Y_{13}-Y_{11}+Y_{22}-Y_{33}.$
+The constraint $X_{ij}\in\operatorname{conv}(SO(3))$ has been shown to be equivalent to a semidefinite constraint [41, 40].
+A $3\times 3$ matrix $Y$ is in $\operatorname{conv}(SO(3))$ if and only if $\mathcal{A}(Y)+{\mathtt{I}}\succeq 0$, where $\mathcal{A}(Y)=$
+$\tiny\+pmatrix[colsep=0pt,rowsep=0.5pt]-Y_{11}-Y_{22}+Y_{33}&Y_{13}+Y_{31}&Y_{12}-Y_{21}&Y_{23}+Y_{32}\\ Y_{13}+Y_{31}&Y_{11}-Y_{22}-Y_{33}&Y_{23}-Y_{32}&Y_{12}+Y_{21}\\ Y_{12}-Y_{21}&Y_{23}-Y_{32}&Y_{11}+Y_{22}+Y_{33}&Y_{31}-Y_{13}\\ Y_{23}+Y_{32}&Y_{12}+Y_{21}&Y_{31}-Y_{13}&-Y_{11}+Y_{22}-Y_{33}.$
 In contrast to the relaxation in (SDP-O(3)-iso), the inclusion of the convex hull constraints can rule out solutions with incorrect determinants, as we have shown previously.
@@
 5 Experiments
-We implemented the SDP program using the conic splitting solver [31] with the JuMP [24] wrapper in Julia [^4] Code can be found at: https://ylochman.github.io/anisotropic-ra.
+We implemented the SDP program using the conic splitting solver [34] with the JuMP [27] wrapper in Julia [^6] The code is available at: https://github.com/ylochman/anisotropic-ra.
 The cost matrix in (SDP-cSO(3)) is down-scaled by the average (across the observed relative poses in the scene) of the largest eigenvalue of the computed Hessian matrix, which proved beneficial empirically.
 The absolute and relative feasibilities are set to $10^{-5}$ and $10^{-6}$, respectively, and infeasibility tolerance is $10^{-8}$.
-The number of iterations is limited to 500 000500000500\,000500 000.
+The number of iterations is limited to 500 000500\,000.
 In practice, the methods converge in far fewer iterations to meet the stopping criterion.
@@
 Scatter plots represent all instances, and the solid lines are the respective medians.
-We compare [^5] Complementing results on both synthetic and real data can be found in the supplementary material.
+We compare [^7] Complementing results on synthetic and real data can be found in the supplementary material.
+These also include comparison to the anisotropic extension of the spectral approach [2].
 three approaches: 1) SDP-O(3)-iso is the standard rotation averaging approach that uses the regular chordal distance and ignores the $\operatorname{conv}(SO(3))$ constraints, 2) SDP-O(3)-aniso uses the anisotropic objective function but ignores the $\operatorname{conv}(SO(3))$ constraints, and 3) SDP-cSO(3) is the proposed method that uses both the anisotropic objective and the $\operatorname{conv}(SO(3))$ constraints.
@@
 For each instance of the problem, we then draw $\Delta\bm{\omega}_{ij}$ from $\mathcal{N}(0,H_{ij}^{-1})$ and compute noisy relative rotations $\tilde{R}_{ij}=e^{[\Delta\bm{\omega}_{ij}]_{\times}}R_{i}R_{j}^{\top}$.
-In Figure 3 we evaluate the effect of using the proposed anisotropic error measurements versus regular chordal distances together with the proposed relaxation versus the standard one.
-We vary the number of cameras from $2$ to $100$, and the proportion of missing data from $0\%$ to $90\%$.
-For each configuration pair, we run $100$ instances of the problem.
-In the top row of Figure 3, we report the distance $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$.
-As expected, using the proposed method (i.e., both new objective and new relaxation) consistently leads to lower errors.
-We record the time-to-solution of the solver, shown in the bottom row of Figure 3.
-In this synthetic setting, the proposed relaxation is more efficient.
-Our intuition is that the introduction of local constraints—$X_{ij}\in\text{conv}(SO(3))$—reduces the oscillations in the employed first-order SDP solver and therefore leads to faster time-to-solution for optimizing both isotropic and anisotropic costs with the proposed constraints.
+[…]
 5.2 Real experiments
-[…]
-We report the obtained rank for each method and the distance $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$, where the ground truth $\{R_{i}\}_{i}$ is from the final reconstruction obtained using the pipeline of [29].
-When SDP-O(3)-aniso does not give a solution $X^{*}$ of rank 3, we compute the closest rank 3 approximation to $X^{*}$ by setting all but the three largest eigenvalues to 00.
+We present the results of applying the proposed framework to a number of public structure from motion datasets  [^8] Available at:
+https://www.maths.lth.se/matematiklth/personal/calle/dataset/dataset.html
+from [17, 32].
+For every pair of images (with 5 or more point matches), we use RANSAC and the 5 point solver [31] to generate an initial solution to the pairwise geometry which is refined, using bundle adjustment [45], into a local minimum of the sum-of-squared reprojection errors.
+Here we fix gauge freedom by assuming that the first camera is fixed, the second camera is at distance $1$ from the first, and represent 3D points with homogeneous coordinates with norm $1$.
+We then compute the second order Taylor approximation of the objective function and marginalize over the position of the second camera and 3D points.
+This gives us a quadratic function, in the axis-angle variable that represents the relative rotation, as described in Section 4.1 and Figure 1, that we use for the anisotropic objective.
+In Table 1 we report the obtained rank for each method and the distance $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$, where the ground truth $\{R_{i}\}_{i}$ is from the final reconstruction obtained using the pipeline of [32].
+When SDP-O(3)-aniso does not give a solution $X^{*}$ of rank 3, we compute the closest rank 3 approximation to $X^{*}$ by setting all but the three largest eigenvalues to 0.
 We then compute a factorization $X^{*}=VV^{\top}$, where $V$ has $3$ columns.
@@
 Dataset Method $\operatorname{rank}(X^{*})$ $\sqrt{\sum_{i}\|R_{i}\!-\!R_{i}^{*}\|_{F}^{2}}$ Runtime, s
-LU Sphinx (70 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.0944 2
+LU Sphinx 70 cameras 85% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.0944 2
 SDP-O(3)-aniso 7 18.6037 460
 SDP-cSO(3) 3 0.0740 5
-Round Church (92 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1399 6
+Round Church 92 cameras 98% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.1399 6
 SDP-O(3)-aniso 6 26.3808 632
 SDP-cSO(3) 3 0.1267 55
-UWO (114 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.3142 14
+UWO 114 cameras 77% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.3142 14
 SDP-O(3)-aniso 6 22.6873 1929
 SDP-cSO(3) 3 0.2274 7
-Tsar Nikolai I (89 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1170 7
+Tsar Nikolai I 89 cameras 87% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.1170 7
 SDP-O(3)-aniso 6 26.8944 1245
 SDP-cSO(3) 3 0.0534 5
-Vercingetorix (69 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.3146 2
+Vercingetorix 69 cameras 77% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.3146 2
 SDP-O(3)-aniso 6 14.8244 242
 SDP-cSO(3) 3 0.2910 4
-Eglise Du Dome (69 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.0546 4
+Eglise Du Dome 69 cameras 98% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.0546 4
 SDP-O(3)-aniso 6 20.1954 1840
 SDP-cSO(3) 3 0.0487 5
-King’s College (77 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1656 4
+King’s College 77 cameras 100% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.1656 4
 SDP-O(3)-aniso 6 17.6508 681
 SDP-cSO(3) 3 0.0796 83
-Kronan (131 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.2149 18
+Kronan 131 cameras 98% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.2149 18
 SDP-O(3)-aniso 6 22.6035 2997
 SDP-cSO(3) 3 0.3892 201
-Alcatraz (133 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.1763 19
+Alcatraz 133 cameras 97% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.1763 19
 SDP-O(3)-aniso 6 22.7623 1931
 SDP-cSO(3) 3 0.1268 107
-Museum Barcelona (133 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.2255 22
+Mus. Barcelona 133 cameras 97% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.2255 22
 SDP-O(3)-aniso 7 30.3428 669
 SDP-cSO(3) 3 0.1310 27
-Temple Singapore (157 cameras) [Uncaptioned image] SDP-O(3)-iso 3 0.2646 31
+Temple Singapore 157 cameras 97% indef. [Uncaptioned image] SDP-O(3)-iso 3 0.2646 31
 SDP-O(3)-aniso 6 26.0824 2313
@@
 SDP-O(3)-iso uses the regular chordal distance and ignores the $\operatorname{conv}(SO(3))$ constraints, SDP-O(3)-aniso uses the proposed anisotropic objective but ignores the $\operatorname{conv}(SO(3))$ constraints, SDP-cSO(3) is the proposed approach that uses both the anisotropic objective and the $\operatorname{conv}(SO(3))$ constraints.
+The first column provides with the information about the dataset: the number of cameras and the percentage of indefinite matrices $M_{ij}$.
 6 Conclusions
-In this work, we propose a way to incorporate anisotropic costs in a certifiably optimal rotation averaging framework.
-We demonstrate how traditional solvers designed to optimize chordal distances fail to optimize the new objective due to ignoring the determinant constraint.
-Our new SDP formulation enforces $\operatorname{conv}(SO(3))$ and our empirical evaluation shows that this relaxation is able to recover global minima.
+In this work, we propose to incorporate anisotropic costs in a certifiably optimal rotation averaging framework.
+We demonstrate how existing solvers designed for chordal distances fail to provide sensible solutions for the new objective as these ignore the determinant constraint.
+Our new SDP formulation enforces $\operatorname{conv}(SO(3))$, and our empirical evaluation shows that this relaxation is able to recover global minima.
 We evaluated the new approach on both synthetic and real datasets and obtained more accurate reconstructions than isotropic methods in most cases, confirming recent observations based on local optimization.
-We believe that this work will provide a good entry for research into anisotropic rotation averaging, a subarea of rotation averaging that has mostly been set aside because of the focus on isotropic cases thus far.
-This work largely focuses on the modeling side and understanding the shortcomings of existing relaxations.
+We believe that our work serves as a good entry point for further investigating the anisotropic rotation averaging setting.
+This work focuses mainly on modeling aspects and understanding the shortcomings of existing relaxations.
 We aim to leverage standard SDP solvers; however, the increased strength of our SDP formulation often comes at the cost of more expensive optimization.
@@
 Appendix A First-Order Uncertainty Propagation
-We leverage Laplace’s approximation, which allows us to identify the Hessian of a non-linear least-squares minimization task as the precision matrix of the posterior $p(\tilde{\bm{\omega}}_{ij}|\text{image $i$ and $j$})$ (up to a positive scale).
-In the following we apply first-order uncertainty propagation to move from the minimal axis-angle parametrization $\bm{\omega}\in\mathbb{R}^{3}$ to a direct parametrization of rotation matrices $R_{ij}\in\mathbb{R}^{3\times 3}$.
+[…]
 For $\Delta\bm{\omega}\approx\mathbf{0}$ we have $\exp([\Delta\bm{\omega}]_{\times})\approx{\mathtt{I}}+[\Delta\bm{\omega}]_{\times}$.
@@
 Therefore the Gaussian transformed by the mapping $\Delta\bm{\omega}\to\operatorname{vec}({\mathtt{I}}+[\Delta\bm{\omega}]_{\times})$ has covariance matrix
-$\mathtt{M}^{-1}=\Sigma_{[\Delta\bm{\omega}]_{\times}}=\mathtt{J}\Sigma_{\Delta\bm{\omega}}\mathtt{J}^{\top}=\mathtt{J}\mathtt{H}^{-1}\mathtt{J}^{\top}.$ (A.1)
+$\Sigma_{[\Delta\bm{\omega}]_{\times}}=\mathtt{J}\Sigma_{\Delta\bm{\omega}}\mathtt{J}^{\top}=\mathtt{J}\mathtt{H}^{-1}\mathtt{J}^{\top}.$ (A.1)
 Since $\mathtt{J}$ has rank 3, the above relation only fixes 6 out of the 45 degrees of freedom in $\Sigma_{[\Delta\bm{\omega}]_{\times}}$, hence there are 39 d.o.f.
@@
 $\begin{split}&\operatorname{vec}(R-\tilde{R})^{\top}\mathtt{M}\operatorname{vec}(R-\tilde{R})\\ &\doteq\operatorname{vec}(R)^{\top}\mathtt{M}\operatorname{vec}(R)-2\operatorname{vec}(\tilde{R})^{\top}\mathtt{M}\operatorname{vec}(R)\end{split}$ (A.2)
-is linear in $R$ and matches $\Delta\bm{\omega}^{\top}\mathtt{H}\Delta\bm{\omega}$ to first order. We consider
+is linear in $R$ and matches $\Delta\bm{\omega}^{\top}\mathtt{H}\Delta\bm{\omega}$ to first order for a suitable matrix $\mathtt{M}$ to be determined. We consider
 $\begin{split}\Delta\bm{\omega}^{\top}\mathtt{H}\Delta\bm{\omega}&\stackrel{{\scriptstyle!}}{{=}}\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}\mathtt{M}[\Delta\bm{\omega}]_{\times})\\ {}&=\operatorname{vec}([\Delta\bm{\omega}]_{\times})^{\top}\operatorname{vec}(\mathtt{M}[\Delta\bm{\omega}]_{\times})\\ {}&=(\mathtt{J}\Delta\bm{\omega})^{\top}\operatorname{vec}(\mathtt{M}[\Delta\bm{\omega}]_{\times})\\ {}&=\Delta\bm{\omega}^{\top}\mathtt{J}^{\top}({\mathtt{I}}\otimes\mathtt{M})\operatorname{vec}([\Delta\bm{\omega}]_{\times})\\ {}&=\Delta\bm{\omega}^{\top}\mathtt{J}^{\top}({\mathtt{I}}\otimes\mathtt{M})\mathtt{J}\Delta\bm{\omega}\end{split}$ (A.3)
@@
 $\mathtt{H}=\operatorname{tr}(\mathtt{M}){\mathtt{I}}-\mathtt{H}$ (A.6)
-in accordance with Section 4.1. We relate $[\Delta\bm{\omega}]_{\times}$ and $R$ via $[\Delta\bm{\omega}]_{\times}\approx R\tilde{R}^{\top}-{\mathtt{I}}$ and therefore obtain
+in accordance with Section 4 of the main paper. We relate $\Delta\bm{\omega}]_{\times}$ and $R$ via $\Delta\bm{\omega}]_{\times}\approx R\tilde{R}^{\top}-{\mathtt{I}}$ and therefore obtain
 $\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}$ $\mathtt{M}[\Delta\bm{\omega}]_{\times})\approx\operatorname{tr}\!\left((R\tilde{R}^{\top}-{\mathtt{I}})^{\top}\mathtt{M}(R\tilde{R}^{\top}-{\mathtt{I}})\right)$
@@
 which is equivalent to the cost (9) in the main text.
-A seemingly different solution $\mathtt{M}^{\prime}$ can be obtained by factoring $\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}\mathtt{M}[\Delta\bm{\omega}]_{\times})$ differently,
+Since $\Delta\bm{\omega}]_{\times}$ is (skew-)symmetric, a seemingly different solution $\mathtt{M}^{\prime}$ can be obtained by factoring $\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}\mathtt{M}[\Delta\bm{\omega}]_{\times})$ differently,
 $\begin{split}\Delta\bm{\omega}^{\top}\mathtt{H}\Delta\bm{\omega}&\stackrel{{\scriptstyle!}}{{=}}\operatorname{tr}([\Delta\bm{\omega}]_{\times}^{\top}\mathtt{M}^{\prime}[\Delta\bm{\omega}]_{\times})\\ {}&=\operatorname{vec}(\mathtt{M}^{\prime}[\Delta\bm{\omega}]_{\times})^{\top}\operatorname{vec}([\Delta\bm{\omega}]_{\times})\\ {}&=\operatorname{vec}(\mathtt{M}^{\prime}[\Delta\bm{\omega}]_{\times})^{\top}\mathtt{J}\Delta\bm{\omega}\\ {}&=\left((\mathtt{M}^{\prime}\otimes{\mathtt{I}})\operatorname{vec}([\Delta\bm{\omega}]_{\times})\right)^{\top}\mathtt{J}\Delta\bm{\omega}\\ {}&=\Delta\bm{\omega}^{\top}\mathtt{J}^{\top}(\mathtt{M}^{\prime}\otimes{\mathtt{I}})\mathtt{J}\Delta\bm{\omega}\end{split}$ (A.8)
@@
 There are different ways of parametrizing rotations.
-In the main paper we use the $R=e^{[\Delta\bm{\omega}]_{\times}}\tilde{R}$ which lead to the objective of the form $-\langle M\tilde{R},R\rangle$.
+In the main paper we use the $R=e^{[\Delta\bm{\omega}]_{\times}}\tilde{R}$ which lead to the objective of the form $\langle M\tilde{R},R\rangle$.
 In this section, we investigate the effects of switching to other possible parameterizations.
@@
 $-\langle\tilde{R}^{\alpha}M\tilde{R}^{(1-\alpha)},R\rangle.$ (B.12)
-The natural choices for $\alpha$ are 00, $1/2$ and $1$.
+The natural choices for $\alpha$ are 0, $1/2$ and $1$.
 Setting $\alpha=0$ yields the formulation employed in the main paper.
@@
 In our synthetic experiments the best solution is often obtained with $\alpha=1/2$, however, this setting does not outperform the proposed formulation overall.
-Appendix C Experimental Details
+Appendix C Spectral method for anisotropic costs
+We recall that our main objective is
+$\sum_{i,j}\left\langle M_{ij}\tilde{R}_{ij},R_{j}R_{i}^{\top}\right\rangle.$ (C.13)
+Let us assume the noisefree setting, i.e. $\tilde{R}_{ij}=R_{i}R_{j}^{\top}$. The cost matrix $\mathtt{N}$ is then given by
+$\mathtt{N}=\begin{pmatrix}&\vdots&\\ \cdots&M_{ij}R_{j}R_{i}^{\top}&\cdots\\ &\vdots&\end{pmatrix}$. (C.14)
+Consequently,
+$\begin{split}\mathtt{N}\mathtt{R}&=\begin{pmatrix}\vdots\\ \sum_{i}M_{ij}R_{j}R_{i}^{\top}R_{i}\\ \vdots\end{pmatrix}=\begin{pmatrix}\vdots\\ \sum_{i}M_{ij}R_{j}\\ \vdots\end{pmatrix}\\ &=\underbrace{\begin{pmatrix}\ddots&&\\ &\sum_{i}M_{ij}&\\ &&\ddots\end{pmatrix}}_{=:\mathtt{D}}\mathtt{R}.\end{split}$ (C.15)
+Hence, $\mathtt{N}\mathtt{R}=\mathtt{D}\mathtt{R}$ or
+$\mathtt{D}^{-1}\mathtt{N}\mathtt{R}=\mathtt{R},$ (C.16)
+and $\mathtt{R}$ can be extracted as the eigenspace corresponding to the eigenvalue of 1.
+Since $\mathtt{N}$ does not necessarily have rank-3 (even in the noise-free case), we obtain $\mathtt{R}$ as the right singular vectors corresponding to the three smallest singular values of $\mathtt{N}-{\mathtt{I}}$.
+We include results for this method in Table D.1 introduced in the next section.
+Appendix D Experimental Details
 We provide complementing results on synthetic and real datasets below.
-C.1 Synthetic experiments
-Figure C.1: The configuration of the cameras (top left).
+D.1 Synthetic experiments
+Figure D.1: The configuration of the cameras (top left).
 Rotation error wrt.
 ground truth $\sqrt{\sum_{i}\|R_{i}-R_{i}^{*}\|_{F}^{2}}$ (top right, bottom) for the increasing noise scale around one of the axes.
-We set up a synthetic graph with three cameras as shown in Figure C.1.
+We set up a synthetic graph with three cameras as shown in Figure D.1.
 The estimated relative rotations corresponding to the black dashed edges are “certain”, i.e.
@@
 The estimated relative rotation of the gray edge has varying uncertainty around one of the three axes, e.g., for the x- (red) axis, the noise covariance is $\operatorname{diag}(\sigma,\varepsilon,\varepsilon)$, where $\sigma$ varies from $0.01$ to $0.3$.
-The effects on the error are shown in Figure C.1.
+The effects on the error are shown in Figure D.1.
 The proposed method relies on certain relative rotations and gives an accurate solution, while the standard isotropic approach is negatively affected by the single noisy relative rotation.
@@
 (f) $p=1.0$
-Figure C.2: Histograms of rotation errors wrt.
+Figure D.2: Histograms of rotation errors wrt.
 ground truth (rows 1,3,5).
@@
 We also present complementing results of the synthetic experiments of Sec 5.1 in the paper.
-Figure C.2 shows the rotation error histograms and the dependency of the runtime on the number of cameras for the other fractions of the observed relative rotations.
+Figure D.2 shows the rotation error histograms and the dependency of the runtime on the number of cameras for the other fractions of the observed relative rotations.
 These results are in line with the analysis provided in the paper — using the proposed objective leads to lower errors and using the proposed constraints speeds up the employed SDP solver.
-C.2 Real experiments
+D.2 Real experiments
 To account for the estimated uncertainties in the evaluation, we compared the methods using the Mahalanobis distance between the axis-angle vectors of the estimated and ground truth rotations.
@@
 Let $\Delta\bm{\omega}_{i}^{-}=\bm{\omega}_{i}-\bm{\omega}_{i}^{*}$ and $\Delta\bm{\omega}_{i}^{+}=\bm{\omega}_{i}+\bm{\omega}_{i}^{*}$, then the Mahalanobis error is
-$\sqrt{\sum_{i}\min\{\Delta\bm{\omega}_{i}^{-\top}H_{i}\Delta\bm{\omega}_{i}^{-},\Delta\bm{\omega}_{i}^{+\top}H_{i}\Delta\bm{\omega}_{i}^{+}\}},$ (C.13)
+$\sqrt{\sum_{i}\min\{\Delta\bm{\omega}_{i}^{-\top}H_{i}\Delta\bm{\omega}_{i}^{-},\Delta\bm{\omega}_{i}^{+\top}H_{i}\Delta\bm{\omega}_{i}^{+}\}},$ (D.17)
 where the minimization is done to account for the sign ambiguity in the axis-angle representation.
 We also present the RMS angular errors for a better geometric interpretation.
-As shown in Table C.1, in many cases, the proposed method leads to a much lower error.
+As shown in Table D.1, in many cases, the proposed method leads to a much lower error.
 Dataset Method Mahal. err. Angl. err.
-LU Sphinx SDP-O(3)-iso 0.3879 0.46
-SDP-cSO(3) 0.2067 0.36
-Round Church SDP-O(3)-iso 0.6308 0.59
-SDP-cSO(3) 0.3683 0.54
-UWO SDP-O(3)-iso 1.4814 1.19
-SDP-cSO(3) 0.7267 0.86
-Tsar Nikolai I SDP-O(3)-iso 0.6871 0.48
-SDP-cSO(3) 0.1878 0.22
-Vercingetorix SDP-O(3)-iso 0.4312 1.53
-SDP-cSO(3) 0.4228 1.42
-Eglise Du Dome SDP-O(3)-iso 0.2239 0.24
-SDP-cSO(3) 0.1882 0.21
-King’s College SDP-O(3)-iso 0.2288 0.76
-SDP-cSO(3) 0.1303 0.37
-Kronan SDP-O(3)-iso 0.7380 0.76
-SDP-cSO(3) 1.1114 1.38
-Alcatraz SDP-O(3)-iso 1.3329 0.62
-SDP-cSO(3) 1.0110 0.45
-Museum Barcelona SDP-O(3)-iso 2.7101 0.79
-SDP-cSO(3) 1.2164 0.46
-Temple Singapore SDP-O(3)-iso 2.4197 0.86
-SDP-cSO(3) 1.0760 0.55
-Table C.1: Mahalanobis distance between the axis-angle vectors of the estimated and ground truth rotations and RMS angular errors (degrees) evaluated on the real datasets.
+LU Sphinx SDP-O(3)-iso 0.388 0.46
+Spectral 0.420 1.17
+SDP-cSO(3) 0.207 0.36
+Round Church SDP-O(3)-iso 0.631 0.59
+Spectral 0.437 1.20
+SDP-cSO(3) 0.368 0.54
+UWO SDP-O(3)-iso 1.481 1.19
+Spectral 6.125 7.07
+SDP-cSO(3) 0.727 0.86
+Tsar Nikolai I SDP-O(3)-iso 0.687 0.48
+Spectral 0.344 0.71
+SDP-cSO(3) 0.188 0.22
+Vercingetorix SDP-O(3)-iso 0.431 1.53
+Spectral 30.970 86.94
+SDP-cSO(3) 0.423 1.42
+Eglise Du Dome SDP-O(3)-iso 0.224 0.24
+Spectral 0.119 0.22
+SDP-cSO(3) 0.188 0.21
+King’s College SDP-O(3)-iso 0.229 0.76
+Spectral 0.251 1.00
+SDP-cSO(3) 0.130 0.37
+Kronan SDP-O(3)-iso 0.738 0.76
+Spectral 2.622 4.36
+SDP-cSO(3) 1.111 1.38
+Alcatraz SDP-O(3)-iso 1.333 0.62
+Spectral 0.667 0.80
+SDP-cSO(3) 1.011 0.45
+Museum Barcelona SDP-O(3)-iso 2.710 0.79
+Spectral 16.588 7.35
+SDP-cSO(3) 1.216 0.46
+Temple Singapore SDP-O(3)-iso 2.420 0.86
+Spectral 0.719 0.46
+SDP-cSO(3) 1.076 0.55
+Table D.1: Mahalanobis distance between the axis-angle vectors of the estimated and ground truth rotations and RMS angular errors (degrees) evaluated on the real datasets.